### v1

# PLD: A Choice-Theoretic List-Wise Knowledge Distillation

Abstract
Knowledge distillation is a model compression technique in which a compact "student" network is trained to replicate the predictive behavior of a larger "teacher" network. In logit-based knowledge distillation, it has become the de facto approach to augment cross-entropy with a distillation term. Typically, this term is either a KL divergence-matching marginal probabilities or a correlation-based loss-capturing intra and inter-class relationships-but in every case it sits as an "add-on" to cross-entropy, with its own weight that must be carefully tuned. In this paper, we adopt a choice-theoretic perspective and recast knowledge distillation under the Plackett-Luce model by interpreting teacher logits as "worth" scores. We introduce Plackett-Luce Distillation (PLD), a weighted list-wise ranking loss in which the teacher model transfers knowledge of its full ranking of classes, weighting each ranked choice by its own confidence. PLD directly optimizes a single "teacher-optimal" ranking-true label first, followed by the remaining classes in descending teacher confidence-yielding a convex, and translation-invariant surrogate that subsumes weighted cross-entropy. Empirically, on standard image-classification benchmarks, PLD improves Top-1 accuracy by an average of +0.42% over (DIST)huang2022knowledge and +1.04% over KDhinton2015distilling in homogeneous settings, and by +0.48% and +1.09% over (DIST) and KD, respectively, in heterogeneous settings.
1 Introduction
Deep neural networks (DNNs) have achieved remarkable success across a wide array of tasks-from image classification and object detection to semantic segmentation and beyond yang2022cloud ; wang2024towards . Accuracy, generalization, and robustness typically improve with increasing depth and parameter count wang2024end , but these gains come at the cost of significantly higher computational and memory demands, limiting deployment in resource-constrained environments such as mobile and embedded devices.
Knowledge distillation (KD) hinton2015distilling addresses this by training a compact “student” network to mimic a large “teacher” network. Classically, KD minimizes a weighted sum of cross-entropy on hard labels and a temperature-scaled Kullback-Leibler (KL) divergence between teacher and student logits. This simple framework enriches the student’s learning signal and has seen widespread adoption for model compression and robustness enhancement gou2021knowledge ; song2022spot .
Despite its popularity, recent studies reveal a paradox: distilling from ever-larger or more accurate teachers can sometimes degrade student performance, particularly under capacity mismatch mirzadeh2020improved ; son2021densely ; huang2022knowledge . A variety of techniques have been proposed to address this issue-teacher-assistant frameworks mirzadeh2020improved ; son2021densely , selective distillation cho2019efficacy ; zhu2022teach , and auxiliary classifiers liang2024neighbor -but these approaches introduce additional architectural components or training stages. An alternative line of work, DIST huang2022knowledge , remains within the classical KD pipeline yet replaces the KL-based distillation term. DIST argues that matching only marginal probabilities via KL divergence fails to preserve the relational structure encoded in the teacher’s outputs. Instead, DIST employs a Pearson correlation-based loss to preserve both inter-class correlations within each prediction and intra-class correlations across examples. Nevertheless, like other distillation methods, DIST still requires a separate cross-entropy term. As Figure 1(a) shows, reducing the weight on this cross-entropy term can improve accuracy, but removing it entirely leads to performance drop. Consequently, a unified objective that subsumes both cross-entropy and distillation-deriving its weights from the teacher’s confidences rather than manual tuning-would be highly desirable.
(a) CE-weight sensitivity for KD and DIST.
(b) KD, DIST, and PLD after 100 vs. 300 epochs.
Figure 1: (a) Varying the CE mixing weight $\alpha$ reveals that KD and DIST have different sensitivities-too much CE hurts both, while a sweet spot near $\alpha\approx 0.1$ maximizes Top-1 accuracy. (b) Under extended training (100 vs. 300 epochs), PLD consistently outperforms both KD and DIST, demonstrating its sustained gains.
To overcome these limitations, we propose a list-wise, choice-theoretic perspective on distillation. We interpret logits as “worth” scores in the classical Plackett-Luce permutation model luce1959individual ; plackett1975analysis , derive a single teacher-optimal ranking-true label first, followed by the teacher’s remaining classes in descending confidence-and impose it on the student via the unsoftened PL likelihood. The PL model is grounded in Luce’s Choice Axiom, which ensures that choice-probability ratios are invariant to irrelevant alternatives, and in Plackett’s extension to full rankings, which yields a coherent distribution over permutations by sequential removal. Because the PL model’s first selection term coincides with cross-entropy on the true label, it inherently subsumes the standard KD cross-entropy loss. However, similar to reddi2021rankdistil ; lan2014position , we observe that proper weighting of each subsequent selection is crucial (see Sec. 5). By weighting step $k$ by the teacher’s softmax mass $\alpha_{k}$, we introduce the Plackett-Luce Distillation (PLD) Loss, a convex, translation-invariant surrogate that unifies cross-entropy, ListMLE xia2008listwise , and P-ListMLE lan2014position as special cases. This choice-theoretic foundation captures the full ranking structure provided by the teacher and enables efficient, gradient-based optimization, eliminating the need for separate cross-entropy terms or manual weight adjustments.
Our main contributions are:
  • We introduce a novel list-wise distillation objective that enforces the teacher-optimal permutation $\pi^{*}$-the unique ordering that places the ground-truth label first and ranks all other classes by descending teacher logits-via the Plackett-Luce likelihood, with each selection step $k$ weighted by the teacher’s softmax mass $\alpha_{k}$, thereby eliminating the need for ad-hoc weight tuning.
  • We show PLD is convex, smoothly differentiable with closed-form gradients, and subsumes CE, ListMLE, and P-ListMLE. Like classical KD and DIST, PLD is simple and efficient, easy to implement (see Appendix C), and requires no architectural modifications.
  • We empirically demonstrate on the image-classification benchmark that PLD consistently outperforms both traditional KD hinton2015distilling and DIST huang2022knowledge : in homogeneous settings, PLD improves Top-1 accuracy by an average of +0.42% over DIST and +1.04% over KD; with a fixed ResNet-50 student distilled from diverse teachers, it achieves gains of +0.48% versus DIST and +1.09% versus KD.
The rest of the paper is organized as follows. Section 2 reviews prior work on KD and ranking losses. Section 3 presents necessary preliminaries. Section 4 derives the PLD loss and analyzes its properties. Section 5 reports empirical results. Finally, Section 6 concludes with a discussion of future directions.
2 Related Work
2.1 Knowledge Distillation and Capacity-Mismatch Remedies
Knowledge distillation hinton2015distilling , inspired by model-bootstrapping techniques zhou2016learnware , trains a compact student to mimic a larger teacher by minimizing a temperature-scaled KL divergence between their outputs logits (softened probabilities) to transfer accuracy and robustness gou2021knowledge while achieving model compression song2022spot . Over time, this simple yet effective framework has become a cornerstone for compressing and enhancing deep neural networks. However, a range of studies have shown that students distilled from very large or highly accurate teachers can underperform, even under adversarially robust settings cho2019efficacy ; park2019relational ; mirzadeh2020improved ; son2021densely ; zhu2021student ; wang2022efficient ; zhu2022teach ; huang2022knowledge ; rao2023parameter ; liang2024neighbor ; yuan2024student ; yin2024adversarial . This counterintuitive degradation is mainly ascribed to a capacity mismatch between teacher and student. To address this, architecture-level innovations introduce intermediate or auxiliary models-Teacher Assistant KD (TAKD) employs a mid-sized assistant to smooth the gap mirzadeh2020improved , Densely Guided KD (DGKD) aggregates multiple assistants for richer supervision son2021densely , Neighbor Self-KD (NSKD) adds auxiliary classifiers within student layers liang2024neighbor , and Student Customized KD (SCKD) gates the distillation loss based on gradient alignment with the student’s own objective zhu2021student . Complementarily, teacher-knowledge regularization techniques refine the source signals: early-stopped teachers often outperform fully converged ones cho2019efficacy , CheckpointKD selects intermediate checkpoints to avoid over-specialization wang2022efficient , undistillable classes are excluded to focus supervision on reliable predictions zhu2022teach , adapter modules control label smoothness rao2023parameter , and Student-friendly KD (SKD) simplifies teacher outputs through a softening-simplifier pipeline yuan2024student .
2.2 Logit-Based Distillation Beyond KL Divergence
This trend is reflected in recent logit-based distillation methods that seek alternatives to KL divergence. Relational KD (RKD) park2019relational transfers mutual relations among data examples via distance and angle-wise losses that penalize structural discrepancies. DIST huang2022knowledge introduced a Pearson correlation coefficient-based loss to explicitly capture the teacher’s intrinsic inter and intra-class relations. Correlation Matching KD (CMKD) niu2024efficient reinterprets inter-class rankings in terms of the decision boundary and employs both Pearson and Spearman correlation losses, dynamically weighted by sample difficulty using a differentiable sorting operation. Additionally, it was observed fan2024revisit that better teacher calibration correlates with improved KD performance under standard methods and therefore was advocated for calibration-insensitive metrics such as ranking-based losses.These methods are computationally efficient-requiring no complex training pipelines or architectural overhead-yet achieve competitive performance. This observation motivates our search for a similarly efficient, calibration-insensitive objective grounded in list-wise ranking theory.
2.3 Ranking-Based and List-Wise Losses for Distillation
Early ranking methods focused on pairwise losses, where each pair of items is classified as correctly or incorrectly ordered-e.g., the ranking SVM (herbrich1999support, ; joachims2002optimizing, ). Listwise approaches, which operate over entire item lists, later gained traction in information retrieval and machine learning (cao2007learning, ; xia2008listwise, ). Many draw on classical statistical ranking models-most notably Luce’s choice model (luce1959individual, ) and the Plackett-Luce permutation distribution (plackett1975analysis, ). However, as the number of items grows, even stochastic top-$k$ extensions like ListNet (luo2015stochastic, ) become impractical, and these losses-designed for direct supervision-do not naturally extend to knowledge distillation.
More recent work applies ranking losses to classification and distillation. For multiclass tasks, listwise losses have been proposed wang2021rank4class ; frydenlund2022language , while distillation under a position-aware binary ranking loss has been studied tang2018ranking ; gao2020understanding . The method in tang2018ranking discards items that the teacher ranks low, whereas lee2019collaborative adapts ranking-based distillation to collaborative filtering. Similarly, RankDistil reddi2021rankdistil preserves the teacher’s top-$k$ ordering by matching the student’s item order to the teacher’s and penalizing lower-ranked items.
Building on ListMLE xia2008listwise , we directly optimize the likelihood of a single teacher-optimal permutation under the Plackett-Luce model. Inspired by RankDistil reddi2021rankdistil and position-aware ListMLE (P-ListMLE) lan2014position , we assign greater weight to top-ranked classes than to those ranked lower. Unlike RankDistil, which is primarily designed for general ranking scenarios and optimizes top-$k$ metrics, our focus is on classification tasks where top-1 accuracy is of primary importance.
3 Preliminaries
We first review the multiclass classification and classical knowledge-distillation framework, then introduce the Plackett-Luce permutation model for rankings.
3.1 Multiclass Classification and Knowledge Distillation
Let $\mathcal{D}=\{(\mathbf{x}^{(n)},y^{(n)})\}_{n=1}^{N}$ be a training set of $N$ examples, where each $\mathbf{x}^{(n)}\in\mathbb{R}^{d}$ and $y^{(n)}\in\{1,\dots,C\}$. A neural network $f_{\theta}:\mathbb{R}^{d}\to\mathbb{R}^{C}$ maps an input $x$ to a logit vector $s\;=\;f_{\theta}(x)\;=\;\bigl{(}s_{1},\dots,s_{C}\bigr{)}\in\mathbb{R}^{C},$ which we refer to as the network’s logits $s\in\mathbb{R}^{C}$. These logits induce a softmax distribution over classes:
$p(y=i\mid s)=\frac{\exp(s_{i})}{\sum_{j=1}^{C}\exp(s_{j})},\quad i=1,\dots,C.$
The standard cross-entropy loss is
$\mathcal{L}_{\mathrm{CE}}(s,y)=-\log p(y\mid s)=-\,s_{y}+\log\!\sum_{j=1}^{C}e^{s_{j}}\,,$
which depends only on the target logit $s_{y}$ and is therefore intransitive: any permutation of the other logits leaves $\mathcal{L}_{\mathrm{CE}}$ unchanged.
In knowledge distillation, a pretrained teacher $f^{T}$ and a student $f^{S}$ produce logits $t=f^{T}(x),\qquad s=f^{S}(x)$.
We soften these via temperature $\tau>0$:
$q^{T}_{i}=\frac{\exp(t_{i}/\tau)}{\sum_{j=1}^{C}\exp(t_{j}/\tau)},\qquad q^{S}_{i}=\frac{\exp(s_{i}/\tau)}{\sum_{j=1}^{C}\exp(s_{j}/\tau)},$
and measure their divergence via a temperature-scaled KL term:
$\mathcal{L}_{\mathrm{KD}}(s,t)=\tau^{2}\,\mathrm{KL}(q^{T}\|q^{S})=\tau^{2}\sum_{i=1}^{C}q^{T}_{i}\log\frac{q^{T}_{i}}{q^{S}_{i}}.$
The student minimizes the combined loss
$\mathcal{L}(s,y)=\alpha\,\mathcal{L}_{\mathrm{CE}}(s,y)+(1-\alpha)\,\mathcal{L}_{\mathrm{KD}}(s,t),$
where $\alpha\in[0,1]$ balances fitting the hard labels against matching the teacher’s full output distribution. Minimizing $\mathcal{L}$ encourages the student both to place the correct class first and to mirror the teacher’s full distribution. In Section 4, we extend this perspective by imposing a full Plackett-Luce ranking over the logits.
3.2 Plackett-Luce (PL) Ranking Model
Let $s\in\mathbb{R}^{C}$ be the logit vector produced by the network. We write $s_{i}$ for its $i$th component and interpret $i\succ j\quad\Longleftrightarrow\quad s_{i}>s_{j}$ as “class $i$ preferred to class $j$.” A full ranking is a permutation $\pi=(\pi_{1},\dots,\pi_{C})\in S_{C}$. The Plackett-Luce model assigns each class $i$ a strictly positive worth $w_{i}=\phi(s_{i})>0,$ where $\phi:\mathbb{R}\to\mathbb{R}_{>0}$ is strictly increasing (we take $\phi(s)=e^{s}$). It defines a probability over permutations by
$P_{\mathrm{PL}}(\pi\mid s)=\prod_{k=1}^{C}\frac{w_{\pi_{k}}}{\sum_{l=k}^{C}w_{\pi_{l}}}=\prod_{k=1}^{C}\frac{\exp(s_{\pi_{k}})}{\sum_{l=k}^{C}\exp(s_{\pi_{l}})}.$
At each step $k$, the class $\pi_{k}$ is chosen from the remaining items in proportion to its worth.
A key property is translation-invariance: adding $\delta$ to all logits (i.e. $s_{i}\mapsto s_{i}+\delta$) multiplies every $w_{i}$ by $e^{\delta}$, which cancels in the softmax-style ratios. By contrast, scaling logits by $a>0$ (i.e. $s_{i}\mapsto a\,s_{i}$) changes the sharpness of the distribution.
Since PL factorizes the full $C!$-way ranking distribution into chained softmax terms, it provides an efficient likelihood for any fixed permutation (xia2008listwise, ). In Section 4 we fix a “teacher-optimal” permutation $\pi^{*}$ and minimize its negative log-likelihood under the student, thereby enforcing both top-1 correctness and richer inter-class ordering information.
4 Method
We derive our PLD objective in two stages: First, we extract a teacher-optimal permutation from the teacher’s logits via the Plackett-Luce model (Sec. 4.1), thereby identifying the core ranking target for distillation. Second, we introduce a confidence-weighted likelihood to obtain the final Plackett-Luce Distillation (PLD) loss (Sec. 4.2).
4.1 Teacher-Optimal Permutation under Plackett-Luce
From the Plackett-Luce ranking viewpoint, a classifier’s logits define a sequential choice process over the $C$ classes. In the first selection step, the model chooses one class out of all $C$ in proportion to $\exp(s_{i})$, exactly matching the softmax probability used in standard cross-entropy. Consequently, minimizing cross-entropy enforces only this first-choice probability-i.e., the probability of picking the correct class first-while remaining agnostic to the ordering of the other $C-1$ classes. In contrast, the full Plackett-Luce model assigns distinct probabilities to all $C!$ permutations by chaining these softmax-style choices at each subsequent step. Thus, cross-entropy can be seen as a relaxed Plackett-Luce loss that cares only about the top-1 selection.
If we augment cross-entropy with a knowledge-distillation term-whether via KL matching of softened distributions or via correlation-based losses-the resulting objective still places its primary emphasis on the true class (via the cross-entropy term) while the distillation term modulates the student’s remaining logits. Viewed through the Plackett-Luce lens, this corresponds to enforcing the first selection to be the ground-truth label, exactly as in cross-entropy, and then shaping the subsequent choices according to the teacher’s preferences rather than forcing an exact match of the full distribution. Concretely, since Plackett-Luce makes each selection independently-choosing $\pi_{k}$ among the remaining classes in proportion to their worth-we fix the first pick to the true label and let the teacher’s descending-logit ordering guide all later picks. Denoting the teacher’s logit vector by $t=(t_{1},\dots,t_{C})\in\mathbb{R}^{C}$, we refer to the resulting fixed ordering
$\pi^{*}\;=\;\bigl{(}y,\;\mathrm{argsort}(t)\setminus\{y\}\bigr{)}$
as the teacher-optimal permutation, and use $\pi^{*}$ as our sole ranking target.
It was shown in xia2008listwise that one can construct surrogate losses on a single Plackett-Luce permutation-using, for example, likelihood (ListMLE), cosine similarity, or cross-entropy-that are all amenable to gradient-based optimization. Among these, the maximum-likelihood (ListMLE) surrogate stands out for its simplicity and for satisfying desirable properties such as consistency, soundness, continuity, differentiability, and convexity, while also exhibiting strong empirical performance. Accordingly, we adopt the ListMLE approach: we place an empirical one-hot target on the teacher-optimal permutation $\pi^{*}$ and define our unweighted loss as its negative log-likelihood under the student’s logits,
$\mathcal{L}_{\mathrm{unweighted}}(s;\pi^{*})\;=\;-\,\log P_{\mathrm{PL}}\bigl{(}\pi^{*}\mid s\bigr{)}\;=\;-\,\sum_{k=1}^{C}\log\frac{\exp\bigl{(}s_{\pi^{*}_{k}}\bigr{)}}{\sum_{\ell=k}^{C}\exp\bigl{(}s_{\pi^{*}_{\ell}}\bigr{)}}.$
This formulation naturally yields a one-hot target over the $C!$ Plackett-Luce permutations-assigning probability one to $\pi^{*}$ and zero to all others-and serves as the foundation for our weighted distillation objective.
4.2 Confidence-Weighted Likelihood Yielding PLD
In the standard PL model, each selection step is weighted equally-i.e., every position contributes identically to the likelihood. Position-aware ListMLE (P-ListMLE) lan2014position addresses this by introducing a fixed, strictly decreasing weight sequence-e.g. $\alpha_{k}=2^{\,C-k}-1,$
so that errors in top-ranked positions are penalized more heavily than those further down. However, in multiclass classification-where only the top-1 decision ultimately matters-fixed, hand-crafted weight schedules are both hyperparameter-sensitive and fail to reflect the inherently greater importance of the first selection compared to subsequent ones. Observing that a pretrained teacher naturally assigns far greater confidence to the correct class than to the others, we therefore parameterize the weight at step $k$ by the teacher’s softmax probability: $\alpha_{k}=q^{T}_{\pi^{*}_{k}}$
This choice ensures the loss automatically emphasizes the top-1 selection when the teacher is confident yet relaxes its focus when the teacher’s output distribution is more uniform. Furthermore, our data-driven weights recover several well-known ranking surrogates as special cases: setting $\alpha=(1,0,\dots,0)$ recovers standard cross-entropy; choosing a uniform distribution $\alpha=(\tfrac{1}{C},\dots,\tfrac{1}{C})$ yields a ListMLE-like objective; and any fixed, strictly decreasing sequence reproduces the P-ListMLE surrogate.
(a) Top-1 accuracy in the homogeneous distillation setting.
(b) Top-1 accuracy in the heterogeneous distillation setting.
Figure 2: (a) Homogeneous setting: larger teachers and smaller students within the same architecture family. (b) Heterogeneous setting: a fixed ResNet-50 student distilled from diverse teacher architectures.
The resulting Plackett-Luce Distillation (PLD) Loss is defined as
$\mathcal{L}_{\mathrm{PLD}}(s,t;y)\;=\;\sum_{k=1}^{C}q^{T}_{\pi^{*}_{k}}\Bigl{[}-\,s_{\pi^{*}_{k}}+\log\!\sum_{\ell=k}^{C}e^{s_{\pi^{*}_{\ell}}}\Bigr{]},$
thereby providing a unified, confidence-weighted ranking objective for knowledge distillation. Since PLD can be viewed as a teacher-softmax-weighted variant of ListMLE, it inherits the same favorable structural properties. In particular, each summand
$-\,s_{\pi^{*}_{k}}\;+\;\log\!\sum_{\ell=k}^{C}e^{s_{\pi^{*}_{\ell}}}$
is convex in the logits $s$, as it is the sum of an affine function and a log-sum-exp. Because nonnegative weighting by $q^{T}_{\pi^{*}_{k}}$ and summation over $k$ preserve convexity, the full PLD loss remains convex in $s$. This convexity, together with smooth differentiability, ensures efficient, gradient-based optimization. The gradient derivation is provided in Appendix A.
5 Experiments
All experiments are conducted on ImageNet-1K deng2009imagenet following the recipe wightman2021resnet with minimal data augmentation. We train for 100 epochs with an effective batch size of 2048 images (256 per GPU across eight NVIDIA A100 SXM4 80 GB accelerators) using the LAMByou2019large optimizer, an initial learning rate of $5\times 10^{-3}$ decayed by a cosine schedule and linearly warmed up over the first 5 epochs, and weight decay fixed at 0.02. Under this setup, 100 epochs require approximately 6 hours for convolutional backbones and 8 hours for Vision Transformers. To focus on the impact of our loss formulations, we apply only random resized crops, horizontal flips, and per-channel normalization to ImageNet statistics.
For validation we adopt the “A-recipe” from wightman2021resnet . Specifically, we set the test resolution $r=224$ and test crop ratio $\rho=0.95$, then $\text{resize}_{\min}=\bigl{\lceil}r/\rho\bigr{\rceil}\approx 236,$ apply a bicubic resize of the shorter side to $\text{resize}_{\min}$, followed by a center crop of size $r\times r$. Finally, we normalize by the same mean and standard deviation used during training.
5.1 Baseline Configuration and Ablation
We evaluate four losses-CE, KD hinton2015distilling , DIST huang2022knowledge , and PLD-using a ResNet-50 student and ResNet-152 teacher He2015 under identical training. Each objective is a drop-in replacement of the logit-based term and incurs comparable cost, so KD (divergence-based) and DIST (correlation-based) serve as natural baselines against our ranking-based PLD. We sweep each method’s key hyperparameters ($\alpha$ for KD, $\beta,\gamma$ for DIST, $\tau_{T}$ for PLD) to find optimal settings. Figure 1(a) shows the effect of varying the CE mixing weight: gradually reducing $\alpha$ improves both KD and DIST up to an optimal region near $\alpha\approx 0.1$, but fully dropping CE ($\alpha=0$) degrades performance. Table 1 reports results with and without the CE term (full ablation in Appendix B), and Table 2 gives PLD-specific sweeps.
Table 1: Top-1 and Top-5 accuracy for baseline methods (ResNet-50 student, ResNet-152 teacher with Top-1 Acc. 79.61%). “–” indicates not applicable. CE: standard cross-entropy. LS:szegedy2016rethinking cross-entropy with label smoothing ($\epsilon=0.1$). KD: vanilla distillation with mixing weight $\alpha$ and temperature $\tau=2$. DIST: relational distillation with inter-class weight $\beta$, intra-class weight $\gamma$, and temperature $\tau=1$.
Method $\alpha$ $\beta$ $\gamma$ $\tau$ Top-1 Acc.Top-5 Acc.%
Teacher (ResNet-152) 79.61
CE 71.35
LS ($\epsilon=0.1$) 73.92
KD 0.00 2.00 75.92 92.82
0.10 2.00 76.80 93.16
DIST 0.00 0.50 0.50 1.00 76.47 93.19
0.10 0.45 0.45 1.00 76.60 93.30
Table 2: Top-1 and Top-5 accuracy for PLD variants (ResNet-50 student, ResNet-152 teacher with Top-1 Acc. 79.61%). “–” indicates not applicable.
Variant $\tau$ Top-1 Acc.Top-5 Acc.%
PLD 0.50 76.06 92.39
1.00 77.30 93.28
1.50 77.17 92.74
2.00 76.46 92.34
4.00 75.22 91.34
PLD–ListMLE 74.11 90.60
PLD–pListMLE 76.60 93.23
5.1.1 Ablation of PLD
As PLD weights each position by $\alpha_{k}\;=\;\frac{\exp\bigl{(}t_{\pi^{*}_{k}}/\tau_{T}\bigr{)}}{\sum_{j=1}^{C}\exp\bigl{(}t_{\pi^{*}_{j}}/\tau_{T}\bigr{)}}$, we sweep $\tau_{T}\in\{0.5,1.0,1.5,2.0,4.0\}$ to assess sensitivity: lower $\tau_{T}$ sharpens the distribution, while higher $\tau_{T}$ flattens it. We also test two special cases: Uniform (ListMLE): $\alpha_{k}=1/C$. Position-aware (p-ListMLE): $\alpha_{k}=q^{T}_{\pi^{*}_{k}}/\sum_{j}q^{T}_{\pi^{*}_{j}}$. As Table 2 shows, the unsoftened PLD ($\tau_{T}=1$) consistently achieves the highest Top-1/Top-5 accuracies, with uniform ListMLE trailing by 2-3 pp. Overall, the best results occur at $\tau_{T}=1.0$ (no softening) or with only slight perturbations ($1.0\pm\epsilon$), demonstrating that PLD is robust to the exact choice of $\alpha$ distribution and does not require aggressive sharpening or over-softening.
5.2 Key Experiments
To evaluate PLD’s robustness and generality, we conduct three complementary studies using hyperparameters identified in our baseline experiments (KD: $\alpha=0.1,\tau=2$; DIST: $\beta=0.45,\gamma=0.45,\tau=1$; PLD: $\tau_{T}=1$). First, in a homogeneous setting, we pair larger teachers with smaller students within the same architecture family to isolate the impact of model capacity on distillation. Second, in a heterogeneous setup, we fix a single student (ResNet-50) and vary teacher architectures across convolutional and transformer families to test PLD under diverse teacher distributions. Finally, we return to the homogeneous scenario with extended pretraining, pretraining a MobileNet-v4qin2024mobilenetv4 Medium student from a MobileNet-v4 Large teacher for 300 epochs to verify that PLD’s gains persist under longer optimization.
5.2.1 Distillation Across Homogeneous Architectures
As illustrated in Figure 2(a), we evaluate three backbone families-ResNetHe2015 , MobileNet-v4qin2024mobilenetv4 , and Vision Transformersdosovitskiy2020image -in a homogeneous setting. Across convolutional models (ResNet, MobileNet-v4), PLD outperforms KD by 0.5-2.7 pp and DIST by 0.2-0.7 pp. On Vision Transformers (ViT), PLD yields +0.72 pp over DIST and +0.30 pp over KD. Averaged across all pairs, PLD improves Top-1 accuracy by +0.42 pp versus DIST and +1.04 pp versus KD. Table 3 reports the full Top-1 results for KD, DIST, and PLD.
5.2.2 Distillation Across Diverse Teachers
One of the key benefits of logit-based distillation is its ability to leverage teacher models of different architectures, provided their output logits share the same dimensionality as the student. Figure 2(b) shows the heterogeneous setting: we fix the student to ResNet-50 ($\approx 25.56$ M parameters) and distill from a range of convolutional and transformer teachers. Across all teachers, PLD consistently outperforms both KD and DIST: its $\Delta_{\mathrm{DIST}}$ gains range from +0.34 pp (MobileNet-v4 Hybrid) to +0.70 pp (ResNet-152), averaging +0.48 pp, while $\Delta_{\mathrm{KD}}$ gains span +0.49 pp to +1.55 pp, with an average of +1.09 pp. As teacher capacity increases, KD’s benefit declines and DIST shows a modest upward trend; PLD not only follows this trend but further widens the margin over DIST for larger teachers. The largest gap between DIST and PLD appears in the homogeneous ResNet-50←ResNet-152 pairing. Table 4 reports each teacher’s standalone Top-1 accuracy, parameter count, and student Top-1 results under KD, DIST, and PLD, along with $\Delta_{\mathrm{DIST}}$ and $\Delta_{\mathrm{KD}}$.
Table 3: Top-1 accuracy (%) and student model size across student–teacher pairs on ImageNet-1K. “Teacher Acc” is the teacher’s standalone Top-1; $\Delta_{\mathrm{DIST}}$ and $\Delta_{\mathrm{KD}}$ are PLD’s gains over DIST and KD.
Top-1 % Accuracy
Student Teacher Params(M) Teacher KD DIST PLD $\Delta_{\mathrm{DIST}}$ $\Delta_{\mathrm{KD}}$
ViT-Small ViT-Large(304.33M) 22.05 84.80 75.33 74.91 75.63 0.72 0.30
ResNet-50 ResNet-152(60.19M) 25.56 79.61 76.80 76.60 77.30 0.70 0.50
MobileNet-v4
Hybrid-Medium Large (conv) 11.07 80.83 75.47 75.98 76.33 0.35 0.86
Medium (conv) (32.59M) 9.72 80.83 74.86 75.40 75.72 0.32 0.86
Small (conv) 3.77 80.83 67.38 70.05 70.07 0.02 2.69
Table 4: Top-1 % accuracy of ResNet-50 distilled from diverse teachers on ImageNet-1K. “Params” is the teacher’s parameter count in millions. $\Delta_{\mathrm{DIST}}$ and $\Delta_{\mathrm{KD}}$ report PLD’s gains over DIST and KD, respectively.
Top-1 % Accuracy $\Delta$
Teacher Params (M) Teacher KD DIST PLD $\boldsymbol{\Delta_{\mathrm{DIST}}}$ $\boldsymbol{\Delta_{\mathrm{KD}}}$
ResNet-152 60.19 79.61 76.80 76.60 77.30 0.70 0.50
MobileNet-v4 Hybrid Medium 11.07 78.66 76.85 77.00 77.34 0.34 0.49
MobileNet-v4 Conv Large 32.59 80.83 73.98 75.05 75.53 0.48 1.55
ViT-Base/16 86.57 82.07 74.27 75.40 75.80 0.40 1.53
ViT-Large/16 304.33 84.80 75.98 76.86 77.38 0.52 1.40
5.2.3 Consistency under Extended Training
As observed in our homogeneous experiments5.2.1 -where KD sometimes outperforms DIST and vice versa-we sought to confirm whether PLD consistently surpasses both. To this end, we independently trained a MobileNet-v4 Medium student from a MobileNet-v4 Large teacher for 100 and 300 epochs using identical hyperparameters. Table 5 reports Top-1 accuracies on ImageNet-1K, and Figure 1(b) provides a detailed comparison. After 100 epochs, PLD achieves 75.72% (vs. DIST +0.32 pp, KD +0.86 pp), rising to 76.94% at 300 epochs (vs. DIST +0.38 pp, KD +0.80 pp). We observe that the distillation gains with PLD (+1.22 pp) are nearly twice those of standard pretraining (+0.67 pp), and that this scaling is comparable across KD (+1.28 pp) and DIST (+1.16 pp). This demonstrates that PLD not only preserves but slightly widens its distillation advantage under extended training.
Table 5: Top-1 accuracy (%) of the MobileNet-v4 Medium student distilled from a MobileNet-v4 Large teacher after 100 and 300 training epochs. The “$\Delta$” row shows the improvement from 100 to 300 epochs.
Epochs CE KD DIST PLD $\Delta_{\mathrm{DIST}}$ $\Delta_{\mathrm{KD}}$
100 epochs 72.37 74.86 75.40 75.72 0.32 0.86
300 epochs 73.04 76.14 76.56 76.94 0.38 0.80
$\Delta$ (300–100) 0.67 1.28 1.16 1.22
6 Conclusion and Limitation
In this work, we introduced Plackett-Luce Distillation (PLD), a unified choice-theoretic framework for logit-based knowledge distillation. Empirically, PLD consistently outperforms both classical KL-based distillation (KD) and recent correlation-based methods (DIST) across a wide range of student-teacher pairs-homogeneous and heterogeneous architectures alike-and remains robust under extended training. These results demonstrate that transferring the teacher’s structured preferences yields stronger and more stable student performance than marginal- or correlation-based objectives.
While PLD applies broadly to any setting in which student and teacher share the same logit dimensionality, it assumes fully aligned class vocabularies and does not directly support mismatched label sets or incremental class addition. Moreover, PLD requires a sorting operation over the $C$ output logits to extract the teacher-optimal permutation, incurring an $O(C\log C)$ complexity per example; although efficient in distributed implementations, this contrasts with the $O(C)$ cost of standard KD or DIST. Finally, as with other distillation methods, PLD’s benefits scale with teacher confidence and may diminish when the teacher’s softmax distribution is nearly uniform.
PLD’s choice-theoretic foundation suggests several promising directions. First, adaptive or curriculum-driven weighting schemes could modulate the ranking loss according to sample difficulty. Second, extending PLD to other domains-such as sequence modeling, reinforcement learning, or multitask settings-may yield similar gains. We hope this work inspires continued exploration of ranking-based objectives for principled and effective model compression.
Appendix A Gradient Derivation for PLD Loss
A.1 Definitions
Let $s=(s_{1},\dots,s_{C})\in\mathbb{R}^{C}$ be the student’s logits. Let $\pi^{*}=(\pi^{*}_{1},\dots,\pi^{*}_{C})$ be a given permutation of class indices, referred to as the teacher-optimal permutation. Let $\alpha_{k}$ for $k=1,\dots,C$ be scalar weights, constant with respect to $s$.
The PLD loss function is defined as:
$\mathcal{L}(s)=\sum_{k=1}^{C}\alpha_{k}\left[-s_{\pi^{*}_{k}}+\log\sum_{\ell=k}^{C}\exp(s_{\pi^{*}_{\ell}})\right].$
We can write $\mathcal{L}(s)=\sum_{k=1}^{C}L_{k}(s)$, where
$L_{k}(s)=\alpha_{k}\left[-s_{\pi^{*}_{k}}+\phi_{k}(s)\right],$
and
$\phi_{k}(s)=\log\sum_{\ell=k}^{C}\exp(s_{\pi^{*}_{\ell}}).$
Our goal is to compute the gradient s($\nabla_{s}\mathcal{L}(s)$, whose $i$-th component is $\frac{\partial\mathcal{L}}{\partial s_{i}}$. Due to the linearity of differentiation,
$\frac{\partial\mathcal{L}}{\partial s_{i}}=\sum_{k=1}^{C}\frac{\partial L_{k}}{\partial s_{i}}.$
We will first compute $\frac{\partial L_{k}}{\partial s_{i}}$.
A.2 Derivative of $L_{k}(s)$:
We have $L_{k}(s)=-\alpha_{k}s_{\pi^{*}_{k}}+\alpha_{k}\phi_{k}(s)$. Let’s differentiate each term with respect to $s_{i}$.
1. Derivative of the affine term:
The first term is $-\alpha_{k}s_{\pi^{*}_{k}}$.
$\frac{\partial}{\partial s_{i}}\left(-\alpha_{k}s_{\pi^{*}_{k}}\right)=-\alpha_{k}\frac{\partial s_{\pi^{*}_{k}}}{\partial s_{i}}.$
Since $s_{\pi^{*}_{k}}$ is the component of $s$ at index $\pi^{*}_{k}$, its derivative with respect to $s_{i}$ is 1 if $i=\pi^{*}_{k}$ and 0 otherwise. This can be written using the indicator function $\mathbf{1}\{i=\pi^{*}_{k}\}$.
$\frac{\partial}{\partial s_{i}}\left(-\alpha_{k}s_{\pi^{*}_{k}}\right)=-\alpha_{k}\mathbf{1}\{i=\pi^{*}_{k}\}.$
2. Derivative of the log-sum-exp term:
The second term is $\alpha_{k}\phi_{k}(s)$, where $\phi_{k}(s)=\log\sum_{\ell=k}^{C}\exp(s_{\pi^{*}_{\ell}})$.
$\frac{\partial}{\partial s_{i}}\left(\alpha_{k}\phi_{k}(s)\right)=\alpha_{k}\frac{\partial\phi_{k}}{\partial s_{i}}.$
To compute $\frac{\partial\phi_{k}}{\partial s_{i}}$, let $X_{k}(s)=\sum_{\ell=k}^{C}\exp(s_{\pi^{*}_{\ell}})$. Then $\phi_{k}(s)=\log X_{k}(s)$. Using the chain rule, $\frac{\partial\phi_{k}}{\partial s_{i}}=\frac{1}{X_{k}(s)}\frac{\partial X_{k}(s)}{\partial s_{i}}$. Now, we compute $\frac{\partial X_{k}(s)}{\partial s_{i}}$:
$\frac{\partial X_{k}(s)}{\partial s_{i}}$ $=\frac{\partial}{\partial s_{i}}\left(\sum_{\ell=k}^{C}\exp(s_{\pi^{*}_{\ell}})\right)$
$=\sum_{\ell=k}^{C}\frac{\partial}{\partial s_{i}}(\exp(s_{\pi^{*}_{\ell}}))$
$=\sum_{\ell=k}^{C}\left(\exp(s_{\pi^{*}_{\ell}})\cdot\frac{\partial s_{\pi^{*}_{\ell}}}{\partial s_{i}}\right)$
$=\sum_{\ell=k}^{C}\exp(s_{\pi^{*}_{\ell}})\mathbf{1}\{i=\pi^{*}_{\ell}\}.$
The term $\mathbf{1}\{i=\pi^{*}_{\ell}\}$ is non-zero (equal to 1) only if $i=\pi^{*}_{\ell}$. This occurs if the index $i$ is part of the set of indices $\{\pi^{*}_{k},\pi^{*}_{k+1},\dots,\pi^{*}_{C}\}$. If $i$ is in this set, then exactly one $\ell$ in the sum (from $k$ to $C$) will satisfy $\pi^{*}_{\ell}=i$, and for that specific $\ell$, the term becomes $\exp(s_{i})$. If $i$ is not in this set, all terms are zero. Thus,
$\frac{\partial X_{k}(s)}{\partial s_{i}}=\begin{cases}\exp(s_{i})&\text{if }i\in\{\pi^{*}_{k},\dots,\pi^{*}_{C}\}\\ 0&\text{otherwise}.\end{cases}$
So, $\frac{\partial\phi_{k}}{\partial s_{i}}$ becomes:
$\frac{\partial\phi_{k}}{\partial s_{i}}=\frac{1}{\sum_{\ell=k}^{C}\exp(s_{\pi^{*}_{\ell}})}\cdot\begin{cases}\exp(s_{i})&\text{if }i\in\{\pi^{*}_{k},\dots,\pi^{*}_{C}\}\\ 0&\text{otherwise}\end{cases}.$
Let us define $\sigma_{k}(i)$ as:
$\sigma_{k}(i):=\begin{cases}\dfrac{\exp(s_{i})}{\sum_{\ell=k}^{C}\exp(s_{\pi^{*}_{\ell}})}&\text{if }i\in\{\pi^{*}_{k},\dots,\pi^{*}_{C}\}\\[12.0pt] 0&\text{otherwise}.\end{cases}$
Then, $\frac{\partial\phi_{k}}{\partial s_{i}}=\sigma_{k}(i)$. And the derivative of the second term of $L_{k}(s)$ is:
$\frac{\partial}{\partial s_{i}}\left(\alpha_{k}\phi_{k}(s)\right)=\alpha_{k}\sigma_{k}(i).$
3. Combining derivatives for $L_{k}(s)$.
Now, we combine the derivatives of the two parts of $L_{k}(s)$:
$\frac{\partial L_{k}}{\partial s_{i}}$ $=\frac{\partial}{\partial s_{i}}\left(-\alpha_{k}s_{\pi^{*}_{k}}\right)+\frac{\partial}{\partial s_{i}}\left(\alpha_{k}\phi_{k}(s)\right)$
$=-\alpha_{k}\mathbf{1}\{i=\pi^{*}_{k}\}+\alpha_{k}\sigma_{k}(i)$
$=\alpha_{k}\left[\sigma_{k}(i)-\mathbf{1}\{i=\pi^{*}_{k}\}\right].$
A.3 Final Gradient s($\nabla_{s}\mathcal{L}(s)$:
The $i$-th component of the gradient of the total loss $\mathcal{L}(s)$ is:
$\frac{\partial\mathcal{L}}{\partial s_{i}}$ $=\sum_{k=1}^{C}\frac{\partial L_{k}}{\partial s_{i}}$
$=\sum_{k=1}^{C}\alpha_{k}\left[\sigma_{k}(i)-\mathbf{1}\{i=\pi^{*}_{k}\}\right].$
We can separate this sum into two parts:
$\frac{\partial\mathcal{L}}{\partial s_{i}}=\sum_{k=1}^{C}\alpha_{k}\sigma_{k}(i)-\sum_{k=1}^{C}\alpha_{k}\mathbf{1}\{i=\pi^{*}_{k}\}.$
For the first part of this expression, $\sum_{k=1}^{C}\alpha_{k}\sigma_{k}(i)$, the term $\sigma_{k}(i)$ is non-zero if and only if $i\in\{\pi^{*}_{k},\dots,\pi^{*}_{C}\}$. Therefore, this sum can be written as $\sum_{k\,:\,i\in\{\pi^{*}_{k},\dots,\pi^{*}_{C}\}}\alpha_{k}\frac{\exp(s_{i})}{\sum_{\ell=k}^{C}\exp(s_{\pi^{*}_{\ell}})}$. For the second part, $\sum_{k=1}^{C}\alpha_{k}\mathbf{1}\{i=\pi^{*}_{k}\}$, the indicator function $\mathbf{1}\{i=\pi^{*}_{k}\}$ is equal to 1 if and only if $\pi^{*}_{k}=i$. Since $\pi^{*}$ is a permutation, for any given $i$, there is precisely one value of $k$ for which this condition is met. Consequently, this sum simplifies to $\sum_{k\,:\,\pi^{*}_{k}=i}\alpha_{k}$, which effectively selects the single $\alpha_{k}$ corresponding to the position of $i$ in the permutation $\pi^{*}$.
Combining these, the $i$-th component of the gradient is:
$\frac{\partial\mathcal{L}}{\partial s_{i}}=\sum_{k:\,i\in\{\pi^{*}_{k},\dots,\pi^{*}_{C}\}}\alpha_{k}\frac{\exp(s_{i})}{\sum_{\ell=k}^{C}\exp(s_{\pi^{*}_{\ell}})}-\sum_{k:\,\pi^{*}_{k}=i}\alpha_{k}.$
In vector form, let $e_{\pi^{*}_{k}}$ denote the standard basis vector that is 1 at index $\pi^{*}_{k}$ and 0 elsewhere. Let $\sigma_{k}$ be a vector whose $j$-th component is $\sigma_{k}(j)$. Then the gradient of $L_{k}(s)$ is $\nabla_{s}L_{k}(s)=\alpha_{k}(\sigma_{k}-e_{\pi^{*}_{k}})$. The total gradient is:
$\nabla_{s}\mathcal{L}(s)=\sum_{k=1}^{C}\alpha_{k}(\sigma_{k}-e_{\pi^{*}_{k}}).$
A.4 Convexity and Smoothness:
Each term $L_{k}(s)=\alpha_{k}\left[-s_{\pi^{*}_{k}}+\log\sum_{\ell=k}^{C}\exp(s_{\pi^{*}_{\ell}})\right]$ is a sum of an affine function ($-\alpha_{k}s_{\pi^{*}_{k}}$) and a non-negatively weighted log-sum-exp function ($\alpha_{k}\phi_{k}(s)$). Affine functions are convex. The log-sum-exp function $\phi_{k}(s)$ is known to be convex. Assuming $\alpha_{k}\geq 0$ (which is true if they are derived from softmax probabilities, as in $q^{T}_{\pi^{*}_{k}}$), the term $\alpha_{k}\phi_{k}(s)$ is also convex. The sum of convex functions is convex, so each $L_{k}(s)$ is convex. The total loss $\mathcal{L}(s)=\sum_{k=1}^{C}L_{k}(s)$ is a sum of convex functions, and therefore, $\mathcal{L}(s)$ is convex in $s$.
Furthermore, affine functions are smooth (infinitely differentiable). The log-sum-exp function is also smooth. Since differentiation, non-negative weighting, and summation preserve smoothness (specifically, Lipschitz continuity of the gradient on any bounded domain), the full PLD loss $\mathcal{L}(s)$ has a Lipschitz-continuous gradient on any bounded domain, which is beneficial for gradient-based optimization methods.
Appendix B Full Baseline Ablation
Table 6: Top-1 and Top-5 accuracy for baseline methods (ResNet-50 student, ResNet-152 teacher with Top-1 Acc. 79.61%). “-” indicates not applicable. CE: standard cross-entropy. LS: cross-entropy with label smoothing ($\epsilon=0.1$). KD: vanilla distillation with mixing weight $\alpha$ and temperature $\tau=2$. DIST: relational distillation with inter-class weight $\beta$, intra-class weight $\gamma$, and temperature $\tau=1$.
Method $\alpha$ $\beta$ $\gamma$ $\tau$ Top-1 Acc.Top-5 Acc.%
Oracle Teacher
Teacher (ResNet-152) - - - - 79.61 -
Baselines
CE - - - - 71.35 -
LS ($\epsilon=0.1$) - - - - 73.92 -
Vanilla KD ($\tau=2$)
KD 0.00 - - 2.00 75.92 92.82
0.10 - - 2.00 76.80 93.16
0.20 - - 2.00 76.65 93.11
0.30 - - 2.00 76.30 93.02
0.40 - - 2.00 76.33 93.10
0.50 - - 2.00 76.12 93.12
0.60 - - 2.00 75.94 92.79
0.70 - - 2.00 75.85 92.71
0.80 - - 2.00 75.21 92.39
0.90 - - 2.00 74.28 91.59
DIST ($\tau=1$)
DIST 1.00 2.00 2.00 1.00 75.77 92.67
0.00 0.50 0.50 1.00 76.47 93.19
0.10 0.45 0.45 1.00 76.60 93.30
0.20 0.40 0.40 1.00 75.69 92.63
0.30 0.35 0.35 1.00 74.88 92.18
0.40 0.30 0.30 1.00 73.90 91.35
0.50 0.25 0.25 1.00 72.94 90.95
0.60 0.20 0.20 1.00 72.71 90.58
0.70 0.15 0.15 1.00 71.85 90.16
0.80 0.10 0.10 1.00 71.89 90.16
0.90 0.05 0.05 1.00 71.40 89.88
Appendix C Implementation of the PLD Loss
1import torch
2import torch.nn.functional as F
3
4def create_adjusted_ranking(flat_logits: torch.Tensor,
5 flat_labels: torch.LongTensor
6 ) -> torch.LongTensor:
7 """
8␣␣␣␣Foreachexample,sortlogitsascending,removethetruelabel,
9␣␣␣␣andappenditattheendsoitoccupiesthelast(top-1)position.
10␣␣␣␣Returnsa[N,V]LongTensorofclassindicesperrank.
11␣␣␣␣"""
12 _, sorted_idx = torch.sort(flat_logits, dim=-1, descending=False)#[N,V]
13 mask = sorted_idx != flat_labels.unsqueeze(-1) #[N,V]
14 V = flat_logits.size(-1)
15
16 assert torch.all(mask.sum(dim=-1) == V - 1), \
17 "Mustremoveexactlyonetruelabelperrow."
18
19 sorted_excl = sorted_idx[mask].view(-1, V - 1) #[N,V-1]
20 return torch.cat([sorted_excl, flat_labels.unsqueeze(-1)], dim=-1)#[N,V]
21
22def plackett_luce_loss(student_logits: torch.Tensor,
23 teacher_logits: torch.Tensor,
24 labels: torch.LongTensor,
25 temperature: float = 1.0
26 ) -> torch.Tensor:
27 """
28␣␣␣␣ComputesthePLDloss:
29␣␣␣␣␣␣L=sum_kalpha_k[logsumexp_k-s_k],
30␣␣␣␣␣␣alpha_k=softmax_teacher[pi*_k].
31␣␣␣␣"""
32 flat_s, flat_t, flat_lbl = prepare_for_classification(
33 student_logits, labels, teacher_logits
34 )
35 ranking = create_adjusted_ranking(flat_t, flat_lbl) # [N, V]
36 s_perm = torch.gather(flat_s, dim=-1, index=ranking) # [N, V]
37 t_perm = torch.gather(flat_t, dim=-1, index=ranking) # [N, V]
38
39 log_cumsum = torch.logcumsumexp(s_perm, dim=-1) # [N, V]
40 per_pos_loss = log_cumsum - s_perm # [N, V]
41 teacher_prob = F.softmax(t_perm / temperature, dim=-1) # [N, V]
42 weighted = per_pos_loss * teacher_prob # [N, V]
43
44 return weighted.sum(dim=-1).mean() # scalar
Listing 1: PyTorch implementation of Plackett-Luce Distillation (PLD) loss.
Appendix D Additional Experimental Setup
For these supplementary experiments, we adopt a homogeneous teacher-student configuration drawn from the MobileNetV4 qin2024mobilenetv4 family. Specifically, the student network is MobileNetV4 Convolutional Small, and the teacher network is MobileNetV4 Hybrid Large. To ensure that any observed differences arise solely from the choice of loss function, we use exactly the same training hyperparameters as in our main experiments, including learning-rate schedules, batch size, weight decay, and the data-augmentation pipeline.
To comprehensively evaluate the relative merits of Plackett-Luce Distillation (PLD) compared to DIST and KD, we conduct experiments across multiple optimizers, various divergence weightings, and several logit-standardization schemes. We further quantify distributional alignment by measuring the KL divergence between student and teacher softmax outputs throughout training. Finally, to gain geometric insight into each loss, we visualize the loss landscapes of PLD, KD, and DIST.
Appendix E Optimizer Ablation Study
In our main experiments, we adopt the Lamb you2019large optimizer as prescribed by the Timm rw2019timm library. To evaluate the robustness of the Plackett-Luce Distillation (PLD) loss under different optimization schemes, we perform an ablation study in which we replace Lamb with three alternatives: AdamW loshchilov2017decoupled , Adan xie2024adan , and AdaBelief zhuang2020adabelief . Table 7 reports the Top-1 accuracy of the MobileNetV4-Small student on ImageNet-1K for each optimizer, under the KD, DIST, and PLD losses. PLD consistently outperforms both KD and DIST, achieving an average Top-1 gain of 0.825% over DIST and 2.21% over KD, with maximum improvements of 0.91% and 2.70%, respectively.
Table 7: Top-1 accuracy (%) of the MobileNetV4-Small student under different optimizers, comparing DIST, KD, and PLD. $\Delta_{\mathrm{DIST}}=\mathrm{PLD}-\mathrm{DIST},\;\Delta_{\mathrm{KD}}=\mathrm{PLD}-\mathrm{KD}$.
Optimizer DIST KD PLD $\Delta_{\mathrm{DIST}}$ $\Delta_{\mathrm{KD}}$
AdaBelief 69.91 68.92 70.80 0.89 1.88
AdamW 69.94 68.89 70.85 0.91 1.96
Adan 70.07 68.52 70.82 0.75 2.30
Lamb (base) 70.41 68.46 71.16 0.75 2.70
Appendix F Logit Standardization Analysis
Logit standardization-centering or normalizing the teacher and student logits before applying the distillation loss-has been proposed sun2024logit to improve optimization stability and distribution alignment. We evaluate four variants on the MobileNetV4-Small student (distilled from MobileNetV4-Hybrid-Large) using the Lamb optimizer. Specifically, we consider:
DIST + std. logits: standardize both teacher and student logits before computing the DIST loss; KD + std. logits: standardize both teacher and student logits before the KL-based KD loss; PLD + std. logits: standardize both teacher and student logits before the PLD loss; PLD + std. teacher logits: standardize only the teacher logits for the PLD softmax weighting.
Table 8 reports Top-1 accuracy on ImageNet-1K for each standardization variant, using the un-standardized PLD baseline of 71.16%. While standardizing logits yields modest gains for both DIST and KD, it does not benefit PLD: applying standardization to either both teacher and student logits or to the teacher logits alone leads to a slight degradation in PLD’s performance.
Table 8: Effect of logit standardization on Top-1 accuracy (%). PLD baseline (no standardization): 71.16%. $\Delta=71.16-\mathrm{Acc}.$
Method Top-1 Acc. (%) $\boldsymbol{\Delta}$
DIST + std. logits 71.12 0.04
KD + std. logits 69.14 2.02
PLD + std. logits 70.66 0.50
PLD + std. teacher logits 70.81 0.35
Appendix G Classification Accuracy Across Divergences
In the main text we adopt the standard knowledge-distillation loss based on the forward Kullback-Leibler divergence:
$\mathcal{L}_{\mathrm{KD}}=\alpha\,\mathcal{L}_{\mathrm{CE}}(z_{s},y)+(1-\alpha)\,D_{\mathrm{KL}}\bigl{(}\operatorname{softmax}(z_{t}/T)\,\|\,\operatorname{softmax}(z_{s}/T)\bigr{)},$
with $\alpha=0.1$ and $T=2$. In addition to this forward KL term, we evaluate two alternatives: the reverse Kullback-Leibler divergence $D_{\mathrm{KL}}\bigl{(}\operatorname{softmax}(z_{s}/T)\,\|\,\operatorname{softmax}(z_{t}/T)\bigr{)}$ and the Jensen-Shannon divergence $D_{\mathrm{JS}}\bigl{(}\operatorname{softmax}(z_{t}/T),\,\operatorname{softmax}(z_{s}/T)\bigr{)}.$ Table 9 reports Top-1 accuracy (%) using each divergence measure. Jensen-Shannon yields a modest gain over forward KL, while reverse KL performs comparably.
Table 9: Top-1 accuracy (%) of vanilla KD under different divergence measures. The PLD baseline (fixed across rows) is 71.16%. $\Delta_{\mathrm{KD}}=\mathrm{PLD}-\mathrm{KD}.$
Divergence KD Top-1 $\boldsymbol{\Delta_{\mathrm{KD}}}$
Forward KL 68.46 2.70
Reverse KL 67.44 3.72
Jensen-Shannon 69.83 1.33
Appendix H Distribution Matching Analysis
We evaluate how well each distillation loss aligns the student’s output distribution with the teacher’s by measuring the KL divergence between their softmax outputs at the end of training. Table 10 reports these KL values for both homogeneous teacher-student pairs (same model family) and heterogeneous pairs (cross-architecture). Lower KL indicates tighter alignment. Although the PLD loss trains the student to respect the teacher’s preference ordering, in the homogeneous setups PLD achieves better alignment than DIST. However, since the KD loss explicitly minimizes a KL term, it yields the lowest divergence overall, outperforming both DIST and PLD.
Table 10: KL divergence of student vs. teacher softmax outputs under DIST, KD, and PLD. Lower is better.
Teacher Student DIST KD PLD
Homogeneous
MobileNetV4-Large MobileNetV4-Medium 0.67 0.55 0.60
MobileNetV4-Large MobileNetV4-Small 0.95 0.85 0.83
MobileNetV4-Large MobileNetV4-Hybrid-Medium 0.64 0.52 0.60
ViT-Large/16 ViT-Base/16 0.47 0.42 0.72
ViT-Large/16 ViT-Small/16 0.55 0.53 0.69
Heterogeneous
MobileNetV4-Large ResNet50 0.71 0.60 0.65
MobileNetV4-Hybrid-Medium ResNet50 1.27 0.27 1.50
ViT-Base/16 ResNet50 0.43 0.39 0.66
ViT-Large/16 ResNet50 0.46 0.44 0.59
Appendix I Loss Landscape Visualization
To gain geometric insight into the optimization landscapes induced by our three distillation losses (DIST, KD, PLD), we plot 3D surfaces and 2D contours over a two-dimensional slice of the student-logit space.
Setup:
Let $t\in\mathbb{R}^{V}$ be a random teacher-logit vector normalized to unit norm, and let $d_{1},d_{2}\in\mathbb{R}^{V}$ be two random orthonormal directions. For a grid of $(\alpha,\beta)\in[-5\|t\|,5\|t\|]^{2}$, we define student logits $s(\alpha,\beta)\;=\;t\;+\;\alpha\,d_{1}\;+\;\beta\,d_{2}$ and compute each loss $L(s(\alpha,\beta),t)$ at every grid point. Figure 3 plots the loss landscapes for DIST huang2022knowledge , KD hinton2015distilling , and PLD (ours). While KD and PLD both induce convex surfaces, DIST dips sharply yet remains effectively planar in the $(\alpha,\beta)$ slice. Moreover, the contour for PLD is more tightly centered around the origin than that of KD.
(a) DIST loss
(b) KD loss
(c) PLD loss
Figure 3: Loss landscapes of three distillation methods: (a) DIST exhibits a sharp dip yet remains effectively planar; (b) KD shows moderate convexity; (c) PLD (ours) exhibits better convexity with contours mostly centered at the origin.
I.1 Temperature Sensitivity of the PLD Loss Surface
To investigate the effect of the teacher-softmax temperature $T$ on the geometry of the PLD loss landscape, we fix the same two-dimensional $(\alpha,\beta)$ slice and compute the PLD surface at four representative temperatures: $T\in\{2.0,1.0,0.5,0.1\}$. Figure 4 shows both the 3D surface and 2D contour plots for each $T$. We observe that reducing $T$ from 2.0 down to 1.0 produces only minor changes, whereas further lowering $T$ below 1.0 flattens the curvature.
(a) $T=2.0$
(b) $T=1.0$
(c) $T=0.5$
(d) $T=0.1$
Figure 4: PLD loss surfaces at different teacher temperatures. (Top row) $T=2.0$ and $T=1.0$; (Bottom row) $T=0.5$ and $T=0.1$. Lowering $T$ below 1.0 flattens convexity.
NeurIPS Paper Checklist
  1. 1.
    Claims
  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
  3. Answer: [Yes]
  4. 2.
    Limitations
  5. Question: Does the paper discuss the limitations of the work performed by the authors?
  6. Answer: [Yes]
  7. 3.
    Theory assumptions and proofs
  8. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
  9. Answer: [Yes]
  10. 4.
    Experimental result reproducibility
  11. Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
  12. Answer: [Yes]
  13. 5.
    Open access to data and code
  14. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
  15. Answer: [Yes]
  16. 6.
    Experimental setting/details
  17. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
  18. Answer: [Yes]
  19. 7.
    Experiment statistical significance
  20. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
  21. Answer: [No]
  22. Justification: We report mean Top-1 accuracies but do not include error bars or multiple runs due to computational constraints.
  23. 8.
    Experiments compute resources
  24. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
  25. Answer: [Yes]
  26. 9.
    Code of ethics
  27. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
  28. Answer: [Yes]
  29. 10.
    Broader impacts
  30. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
  31. Answer: [N/A]
  32. Justification: This work is foundational on model compression and does not present a direct application or risk. Indirectly, more efficient models may reduce energy and hardware costs, but we do not see specific societal harms arising from the method itself.
  33. 11.
    Safeguards
  34. Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
  35. Answer: [N/A]
  36. Justification: No high-risk or dual-use data/models are released in this work
  37. 12.
    Licenses for existing assets
  38. Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
  39. Answer: [Yes]
  40. 13.
    New assets
  41. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
  42. Answer: [N/A]
  43. Justification: We do not release any new datasets or large models with this submission.
  44. 14.
    Crowdsourcing and research with human subjects
  45. Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
  46. Answer: [N/A]
  47. Justification: No human-subject data or crowdsourcing is involved.
  48. 15.
    Institutional review board (IRB) approvals or equivalent for research with human subjects
  49. Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
  50. Answer: [N/A]
  51. Justification: This work does not involve human subjects.
  52. 16.
    Declaration of LLM usage
  53. Question: Does the paper describe the usage of LLMs if it is an important, original, or non-standard component of the core methods in this research? Note that if the LLM is used only for writing, editing, or formatting purposes and does not impact the core methodology, scientific rigorousness, or originality of the research, declaration is not required.
  54. Answer: [N/A]
  55. Justification: No large language models are used in our core methodology.

### v2

@@
 Figure 4: PLD loss surfaces at different teacher temperatures. (Top row) $T=2.0$ and $T=1.0$; (Bottom row) $T=0.5$ and $T=0.1$. Lowering $T$ below 1.0 flattens convexity.
-NeurIPS Paper Checklist
-  1. 1.
-    Claims
-  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
-  3. Answer: [Yes]
-  4. 2.
-    Limitations
-  5. Question: Does the paper discuss the limitations of the work performed by the authors?
-  6. Answer: [Yes]
-  7. 3.
-    Theory assumptions and proofs
-  8. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
-  9. Answer: [Yes]
-  10. 4.
-    Experimental result reproducibility
-  11.
-Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
-  12. Answer: [Yes]
-  13. 5.
-    Open access to data and code
-  14. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
-  15. Answer: [Yes]
-  16. 6.
-    Experimental setting/details
-  17. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
-  18. Answer: [Yes]
-  19. 7.
-    Experiment statistical significance
-  20. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
-  21. Answer: [No]
-  22. Justification: We report mean Top-1 accuracies but do not include error bars or multiple runs due to computational constraints.
-  23. 8.
-    Experiments compute resources
-  24. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
-  25. Answer: [Yes]
-  26. 9.
-    Code of ethics
-  27. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
-  28. Answer: [Yes]
-  29. 10.
-    Broader impacts
-  30. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
-  31. Answer: [N/A]
-  32.
-Justification: This work is foundational on model compression and does not present a direct application or risk.
-Indirectly, more efficient models may reduce energy and hardware costs, but we do not see specific societal harms arising from the method itself.
-  33. 11.
-    Safeguards
-  34.
-Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
-  35. Answer: [N/A]
-  36. Justification: No high-risk or dual-use data/models are released in this work
-  37. 12.
-    Licenses for existing assets
-  38.
-Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
-  39. Answer: [Yes]
-  40. 13.
-    New assets
-  41. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
-  42. Answer: [N/A]
-  43. Justification: We do not release any new datasets or large models with this submission.
-  44. 14.
-    Crowdsourcing and research with human subjects
-  45.
-Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
-  46. Answer: [N/A]
-  47. Justification: No human-subject data or crowdsourcing is involved.
-  48. 15.
-    Institutional review board (IRB) approvals or equivalent for research with human subjects
-  49.
-Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
-  50. Answer: [N/A]
-  51. Justification: This work does not involve human subjects.
-  52. 16.
-    Declaration of LLM usage
-  53.
-Question: Does the paper describe the usage of LLMs if it is an important, original, or non-standard component of the core methods in this research?
-Note that if the LLM is used only for writing, editing, or formatting purposes and does not impact the core methodology, scientific rigorousness, or originality of the research, declaration is not required.
-  54. Answer: [N/A]
-  55. Justification: No large language models are used in our core methodology.

### v3

@@
 In logit-based knowledge distillation, it has become the de facto approach to augment cross-entropy with a distillation term.
-Typically, this term is either a KL divergence-matching marginal probabilities or a correlation-based loss-capturing intra and inter-class relationships-but in every case it sits as an "add-on" to cross-entropy, with its own weight that must be carefully tuned.
-In this paper, we adopt a choice-theoretic perspective and recast knowledge distillation under the Plackett-Luce model by interpreting teacher logits as "worth" scores.
-We introduce Plackett-Luce Distillation (PLD), a weighted list-wise ranking loss in which the teacher model transfers knowledge of its full ranking of classes, weighting each ranked choice by its own confidence.
-PLD directly optimizes a single "teacher-optimal" ranking-true label first, followed by the remaining classes in descending teacher confidence-yielding a convex, and translation-invariant surrogate that subsumes weighted cross-entropy.
-Empirically, on standard image-classification benchmarks, PLD improves Top-1 accuracy by an average of +0.42% over (DIST)huang2022knowledge and +1.04% over KDhinton2015distilling in homogeneous settings, and by +0.48% and +1.09% over (DIST) and KD, respectively, in heterogeneous settings.
+Typically, this term is either a KL divergence that matches marginal probabilities or a correlation-based loss that captures intra- and inter-class relationships.
+In every case, it acts as an additional term to cross-entropy.
+This term has its own weight, which must be carefully tuned.
+In this paper, we adopt a choice-theoretic perspective and recast knowledge distillation under the Plackett–Luce model by interpreting teacher logits as "worth" scores.
+We introduce Plackett-Luce Distillation (PLD), a weighted list-wise ranking loss.
+In PLD, the teacher model transfers knowledge of its full ranking of classes, weighting each ranked choice by its own confidence.
+PLD directly optimizes a single "teacher-optimal" ranking.
+The true label is placed first, followed by the remaining classes in descending teacher confidence.
+This process yields a convex and translation-invariant surrogate that subsumes weighted cross-entropy.
+Empirically, across CIFAR-100, ImageNet-1K, and MS-COCO, PLD achieves consistent gains across diverse architectures and distillation objectives, including divergence-based, correlation-based, and feature-based methods, in both homogeneous and heterogeneous teacher–student pairs.
 1 Introduction
 Deep neural networks (DNNs) have achieved remarkable success across a wide array of tasks-from image classification and object detection to semantic segmentation and beyond yang2022cloud ; wang2024towards .
-Accuracy, generalization, and robustness typically improve with increasing depth and parameter count wang2024end , but these gains come at the cost of significantly higher computational and memory demands, limiting deployment in resource-constrained environments such as mobile and embedded devices.
-Knowledge distillation (KD) hinton2015distilling addresses this by training a compact “student” network to mimic a large “teacher” network.
+Accuracy, generalization, and robustness usually improve as models become deeper and larger wang2024end .
+However, these gains require more computation and memory, which limits deployment on mobile or embedded devices.
+Knowledge distillation (KD) hinton2015distilling addresses this by training a compact "student" network to mimic a large "teacher" network.
 Classically, KD minimizes a weighted sum of cross-entropy on hard labels and a temperature-scaled Kullback-Leibler (KL) divergence between teacher and student logits.
 This simple framework enriches the student’s learning signal and has seen widespread adoption for model compression and robustness enhancement gou2021knowledge ; song2022spot .
-Despite its popularity, recent studies reveal a paradox: distilling from ever-larger or more accurate teachers can sometimes degrade student performance, particularly under capacity mismatch mirzadeh2020improved ; son2021densely ; huang2022knowledge .
-A variety of techniques have been proposed to address this issue-teacher-assistant frameworks mirzadeh2020improved ; son2021densely , selective distillation cho2019efficacy ; zhu2022teach , and auxiliary classifiers liang2024neighbor -but these approaches introduce additional architectural components or training stages.
+Despite its popularity, recent studies reveal a paradox.
+Distilling from larger or more accurate teachers can sometimes degrade the student’s performance, especially when there is a capacity mismatch mirzadeh2020improved ; son2021densely ; huang2022knowledge .
+Several methods have been proposed to address this issue.
+These include teacher-assistant frameworks mirzadeh2020improved ; son2021densely , selective distillation cho2019efficacy ; zhu2022teach , and auxiliary classifiers liang2024neighbor .
+However, these methods often add extra architectural components or training stages.
 An alternative line of work, DIST huang2022knowledge , remains within the classical KD pipeline yet replaces the KL-based distillation term.
@@
 As Figure 1(a) shows, reducing the weight on this cross-entropy term can improve accuracy, but removing it entirely leads to performance drop.
-Consequently, a unified objective that subsumes both cross-entropy and distillation-deriving its weights from the teacher’s confidences rather than manual tuning-would be highly desirable.
+Therefore, a unified objective that combines cross-entropy and distillation is desirable.
+Such an objective should derive its weights from the teacher’s confidence rather than manual tuning.
 (a) CE-weight sensitivity for KD and DIST.
@@
 To overcome these limitations, we propose a list-wise, choice-theoretic perspective on distillation.
-We interpret logits as “worth” scores in the classical Plackett-Luce permutation model luce1959individual ; plackett1975analysis , derive a single teacher-optimal ranking-true label first, followed by the teacher’s remaining classes in descending confidence-and impose it on the student via the unsoftened PL likelihood.
-The PL model is grounded in Luce’s Choice Axiom, which ensures that choice-probability ratios are invariant to irrelevant alternatives, and in Plackett’s extension to full rankings, which yields a coherent distribution over permutations by sequential removal.
-Because the PL model’s first selection term coincides with cross-entropy on the true label, it inherently subsumes the standard KD cross-entropy loss.
+We interpret logits as "worth" scores under the classical Plackett–Luce model luce1959individual ; plackett1975analysis .
+We then derive a single teacher-optimal ranking, where the true label comes first and the remaining classes follow in descending teacher confidence.
+This ranking is imposed on the student using the unsoftened Plackett–Luce likelihood.
+The PL model is based on Luce’s Choice Axiom, which ensures that choice probabilities are invariant to irrelevant alternatives.
+Plackett’s extension generalizes this to full rankings, forming a coherent distribution over permutations through sequential removal.
+The first selection term of the PL model matches the cross-entropy on the true label.
+Thus, it naturally includes the standard KD cross-entropy loss.
 However, similar to reddi2021rankdistil ; lan2014position , we observe that proper weighting of each subsequent selection is crucial (see Sec.
 5).
-By weighting step $k$ by the teacher’s softmax mass $\alpha_{k}$, we introduce the Plackett-Luce Distillation (PLD) Loss, a convex, translation-invariant surrogate that unifies cross-entropy, ListMLE xia2008listwise , and P-ListMLE lan2014position as special cases.
+We weight each step $k$ by the teacher’s softmax mass $\alpha_{k}$.
+This leads to the Plackett–Luce Distillation (PLD) Loss, a convex and translation-invariant surrogate.
+PLD unifies cross-entropy, ListMLE xia2008listwise , and P-ListMLE lan2014position as special cases.
 This choice-theoretic foundation captures the full ranking structure provided by the teacher and enables efficient, gradient-based optimization, eliminating the need for separate cross-entropy terms or manual weight adjustments.
@@
 Like classical KD and DIST, PLD is simple and efficient, easy to implement (see Appendix C), and requires no architectural modifications.
-  • We empirically demonstrate on the image-classification benchmark that PLD consistently outperforms both traditional KD hinton2015distilling and DIST huang2022knowledge : in homogeneous settings, PLD improves Top-1 accuracy by an average of +0.42% over DIST and +1.04% over KD; with a fixed ResNet-50 student distilled from diverse teachers, it achieves gains of +0.48% versus DIST and +1.09% versus KD.
+  • We empirically demonstrate on CIFAR-100, ImageNet-1K, and MS-COCO that PLD achieves consistent gains across diverse architectures and distillation objectives, including divergence-based, correlation-based, and feature-based methods, in both homogeneous and heterogeneous teacher–student pairs.
 The rest of the paper is organized as follows.
@@
 2.1 Knowledge Distillation and Capacity-Mismatch Remedies
-Knowledge distillation hinton2015distilling , inspired by model-bootstrapping techniques zhou2016learnware , trains a compact student to mimic a larger teacher by minimizing a temperature-scaled KL divergence between their outputs logits (softened probabilities) to transfer accuracy and robustness gou2021knowledge while achieving model compression song2022spot .
+Knowledge distillation (KD) hinton2015distilling , inspired by model-bootstrapping techniques zhou2016learnware , trains a compact student to mimic a larger teacher.
+It minimizes a temperature-scaled KL divergence between their output logits (softened probabilities).
+This process transfers both accuracy and robustness gou2021knowledge while enabling model compression song2022spot .
 Over time, this simple yet effective framework has become a cornerstone for compressing and enhancing deep neural networks.
@@
 This counterintuitive degradation is mainly ascribed to a capacity mismatch between teacher and student.
-To address this, architecture-level innovations introduce intermediate or auxiliary models-Teacher Assistant KD (TAKD) employs a mid-sized assistant to smooth the gap mirzadeh2020improved , Densely Guided KD (DGKD) aggregates multiple assistants for richer supervision son2021densely , Neighbor Self-KD (NSKD) adds auxiliary classifiers within student layers liang2024neighbor , and Student Customized KD (SCKD) gates the distillation loss based on gradient alignment with the student’s own objective zhu2021student .
-Complementarily, teacher-knowledge regularization techniques refine the source signals: early-stopped teachers often outperform fully converged ones cho2019efficacy , CheckpointKD selects intermediate checkpoints to avoid over-specialization wang2022efficient , undistillable classes are excluded to focus supervision on reliable predictions zhu2022teach , adapter modules control label smoothness rao2023parameter , and Student-friendly KD (SKD) simplifies teacher outputs through a softening-simplifier pipeline yuan2024student .
+To address this issue, several architecture-level methods introduce intermediate or auxiliary models.
+Teacher Assistant KD (TAKD) employs a mid-sized assistant to bridge the gap mirzadeh2020improved .
+Densely Guided KD (DGKD) aggregates multiple assistants for richer supervision son2021densely .
+Neighbor Self-KD (NSKD) adds auxiliary classifiers within student layers liang2024neighbor .
+Student Customized KD (SCKD) adjusts the distillation loss based on gradient alignment with the student’s objective zhu2021student .
+Teacher-knowledge regularization techniques also refine the teacher’s signal.
+Early-stopped teachers often outperform fully converged ones cho2019efficacy .
+CheckpointKD selects intermediate checkpoints to prevent over-specialization wang2022efficient .
+Some methods exclude undistillable classes to focus supervision on reliable predictions zhu2022teach .
+Adapter modules control label smoothness rao2023parameter , and Student-friendly KD (SKD) simplifies teacher outputs through a softening-simplifier pipeline yuan2024student .
 2.2 Logit-Based Distillation Beyond KL Divergence
@@
 Correlation Matching KD (CMKD) niu2024efficient reinterprets inter-class rankings in terms of the decision boundary and employs both Pearson and Spearman correlation losses, dynamically weighted by sample difficulty using a differentiable sorting operation.
-Additionally, it was observed fan2024revisit that better teacher calibration correlates with improved KD performance under standard methods and therefore was advocated for calibration-insensitive metrics such as ranking-based losses.These methods are computationally efficient-requiring no complex training pipelines or architectural overhead-yet achieve competitive performance.
+It was also observed fan2024revisit that better teacher calibration correlates with improved KD performance under standard methods.
+This finding motivated calibration-insensitive metrics such as ranking-based losses.
+These methods are computationally efficient, require no complex training pipelines or architectural overhead, yet achieve competitive performance.
 This observation motivates our search for a similarly efficient, calibration-insensitive objective grounded in list-wise ranking theory.
 2.3 Ranking-Based and List-Wise Losses for Distillation
-Early ranking methods focused on pairwise losses, where each pair of items is classified as correctly or incorrectly ordered-e.g., the ranking SVM (herbrich1999support, ; joachims2002optimizing, ).
+Early ranking methods focused on pairwise losses.
+In these methods, each pair of items is classified as correctly or incorrectly ordered, as in the ranking SVM (herbrich1999support, ; joachims2002optimizing, ).
 Listwise approaches, which operate over entire item lists, later gained traction in information retrieval and machine learning (cao2007learning, ; xia2008listwise, ).
 Many draw on classical statistical ranking models-most notably Luce’s choice model (luce1959individual, ) and the Plackett-Luce permutation distribution (plackett1975analysis, ).
-However, as the number of items grows, even stochastic top-$k$ extensions like ListNet (luo2015stochastic, ) become impractical, and these losses-designed for direct supervision-do not naturally extend to knowledge distillation.
+As the number of items increases, even stochastic top-$k$ extensions such as ListNet (luo2015stochastic, ) become impractical.
+Moreover, these losses are designed for direct supervision and do not naturally extend to knowledge distillation.
 More recent work applies ranking losses to classification and distillation.
 For multiclass tasks, listwise losses have been proposed wang2021rank4class ; frydenlund2022language , while distillation under a position-aware binary ranking loss has been studied tang2018ranking ; gao2020understanding .
-The method in tang2018ranking discards items that the teacher ranks low, whereas lee2019collaborative adapts ranking-based distillation to collaborative filtering.
+The method in tang2018ranking removes items that the teacher ranks low.
+In contrast, lee2019collaborative adapts ranking-based distillation to collaborative filtering.
 Similarly, RankDistil reddi2021rankdistil preserves the teacher’s top-$k$ ordering by matching the student’s item order to the teacher’s and penalizing lower-ranked items.
@@
 Inspired by RankDistil reddi2021rankdistil and position-aware ListMLE (P-ListMLE) lan2014position , we assign greater weight to top-ranked classes than to those ranked lower.
-Unlike RankDistil, which is primarily designed for general ranking scenarios and optimizes top-$k$ metrics, our focus is on classification tasks where top-1 accuracy is of primary importance.
+Unlike RankDistil, which is designed for general ranking tasks and optimizes top-$k$ metrics, our method focuses on classification tasks.
+In this setting, top-1 accuracy is most important.
 3 Preliminaries
@@
 Let $s\in\mathbb{R}^{C}$ be the logit vector produced by the network.
-We write $s_{i}$ for its $i$th component and interpret $i\succ j\quad\Longleftrightarrow\quad s_{i}>s_{j}$ as “class $i$ preferred to class $j$.” A full ranking is a permutation $\pi=(\pi_{1},\dots,\pi_{C})\in S_{C}$.
+We write $s_{i}$ for its $i$th component and interpret $i\succ j\quad\Longleftrightarrow\quad s_{i}>s_{j}$ as "class $i$ preferred to class jj." A full ranking is a permutation $\pi=(\pi_{1},\dots,\pi_{C})\in S_{C}$.
 The Plackett-Luce model assigns each class $i$ a strictly positive worth $w_{i}=\phi(s_{i})>0,$ where $\phi:\mathbb{R}\to\mathbb{R}_{>0}$ is strictly increasing (we take $\phi(s)=e^{s}$).
@@
 $s_{i}\mapsto a\,s_{i}$) changes the sharpness of the distribution.
-Since PL factorizes the full $C!$-way ranking distribution into chained softmax terms, it provides an efficient likelihood for any fixed permutation (xia2008listwise, ).
-In Section 4 we fix a “teacher-optimal” permutation $\pi^{*}$ and minimize its negative log-likelihood under the student, thereby enforcing both top-1 correctness and richer inter-class ordering information.
+The PL model factorizes the full $C!$-way ranking distribution into a chain of softmax terms.
+This factorization provides an efficient likelihood for any fixed permutation (xia2008listwise, ).
+In Section 4, we fix a "teacher-optimal" permutation $\pi^{*}$ and minimize its negative log-likelihood under the student.
+This approach enforces both top-1 correctness and richer inter-class ordering information.
 4 Method
-We derive our PLD objective in two stages: First, we extract a teacher-optimal permutation from the teacher’s logits via the Plackett-Luce model (Sec.
-4.1), thereby identifying the core ranking target for distillation.
-Second, we introduce a confidence-weighted likelihood to obtain the final Plackett-Luce Distillation (PLD) loss (Sec.
+We derive the PLD objective in two stages.
+First, we extract a teacher-optimal permutation from the teacher’s logits using the Plackett–Luce model (Sec.
+4.1).
+This step identifies the ranking target for distillation.
+Second, we introduce a confidence-weighted likelihood to form the final Plackett–Luce Distillation (PLD) loss (Sec.
 4.2).
 4.1 Teacher-Optimal Permutation under Plackett-Luce
-From the Plackett-Luce ranking viewpoint, a classifier’s logits define a sequential choice process over the $C$ classes.
+From the Plackett-Luce ranking viewpoint, a classifier’s logits define a sequential choice process among the $C$ classes.
 In the first selection step, the model chooses one class out of all $C$ in proportion to $\exp(s_{i})$, exactly matching the softmax probability used in standard cross-entropy.
-Consequently, minimizing cross-entropy enforces only this first-choice probability-i.e., the probability of picking the correct class first-while remaining agnostic to the ordering of the other $C-1$ classes.
-In contrast, the full Plackett-Luce model assigns distinct probabilities to all $C!$ permutations by chaining these softmax-style choices at each subsequent step.
+Consequently, minimizing cross-entropy enforces only the first-choice probability, i.e., the probability of selecting the correct class first.
+It remains agnostic to the ordering of the remaining C1C-1 classes.
+In contrast, the full Plackett–Luce model assigns probabilities to all $C!$ possible permutations.
+It does this by chaining softmax-style selections at each step.
 Thus, cross-entropy can be seen as a relaxed Plackett-Luce loss that cares only about the top-1 selection.
-If we augment cross-entropy with a knowledge-distillation term-whether via KL matching of softened distributions or via correlation-based losses-the resulting objective still places its primary emphasis on the true class (via the cross-entropy term) while the distillation term modulates the student’s remaining logits.
-Viewed through the Plackett-Luce lens, this corresponds to enforcing the first selection to be the ground-truth label, exactly as in cross-entropy, and then shaping the subsequent choices according to the teacher’s preferences rather than forcing an exact match of the full distribution.
+If we augment cross-entropy with a knowledge-distillation term, such as KL matching or correlation-based loss, the objective still focuses mainly on the true class through the cross-entropy component.
+The distillation term only adjusts the remaining logits.
+Viewed through the Plackett–Luce perspective, this process enforces the first selection to be the ground-truth label, as in cross-entropy.
+It then shapes later choices according to the teacher’s preferences instead of matching the entire distribution.
 Concretely, since Plackett-Luce makes each selection independently-choosing $\pi_{k}$ among the remaining classes in proportion to their worth-we fix the first pick to the true label and let the teacher’s descending-logit ordering guide all later picks.
@@
 as the teacher-optimal permutation, and use $\pi^{*}$ as our sole ranking target.
-It was shown in xia2008listwise that one can construct surrogate losses on a single Plackett-Luce permutation-using, for example, likelihood (ListMLE), cosine similarity, or cross-entropy-that are all amenable to gradient-based optimization.
-Among these, the maximum-likelihood (ListMLE) surrogate stands out for its simplicity and for satisfying desirable properties such as consistency, soundness, continuity, differentiability, and convexity, while also exhibiting strong empirical performance.
-Accordingly, we adopt the ListMLE approach: we place an empirical one-hot target on the teacher-optimal permutation $\pi^{*}$ and define our unweighted loss as its negative log-likelihood under the student’s logits,
+As shown in xia2008listwise , one can construct surrogate losses on a single Plackett–Luce permutation.
+Examples include likelihood-based (ListMLE), cosine-similarity, and cross-entropy variants, all suitable for gradient-based optimization.
+Among these, the likelihood (ListMLE) surrogate stands out for its simplicity and for satisfying desirable properties such as consistency, soundness, continuity, differentiability, and convexity, while also exhibiting strong empirical performance.
+We adopt the ListMLE approach.
+We place an empirical one-hot target on the teacher-optimal permutation $\pi^{*}$ and define its negative log-likelihood under the student’s logits as the unweighted loss.
 $\mathcal{L}_{\mathrm{unweighted}}(s;\pi^{*})\;=\;-\,\log P_{\mathrm{PL}}\bigl(\pi^{*}\mid s\bigr)\;=\;-\,\sum_{k=1}^{C}\log\frac{\exp\bigl(s_{\pi^{*}_{k}}\bigr)}{\sum_{\ell=k}^{C}\exp\bigl(s_{\pi^{*}_{\ell}}\bigr)}.$
@@
 In the standard PL model, each selection step is weighted equally-i.e., every position contributes identically to the likelihood.
-Position-aware ListMLE (P-ListMLE) lan2014position addresses this by introducing a fixed, strictly decreasing weight sequence-e.g.
-$\alpha_{k}=2^{\,C-k}-1,$
-so that errors in top-ranked positions are penalized more heavily than those further down.
+Position-aware ListMLE (P-ListMLE) lan2014position addresses this by introducing a fixed, strictly decreasing weight sequence, such as $\alpha_{k}=2^{\,C-k}-1.$
+This scheme penalizes errors at top ranks more than those lower down.
 However, in multiclass classification-where only the top-1 decision ultimately matters-fixed, hand-crafted weight schedules are both hyperparameter-sensitive and fail to reflect the inherently greater importance of the first selection compared to subsequent ones.
@@
 This choice ensures the loss automatically emphasizes the top-1 selection when the teacher is confident yet relaxes its focus when the teacher’s output distribution is more uniform.
-Furthermore, our data-driven weights recover several well-known ranking surrogates as special cases: setting $\alpha=(1,0,\dots,0)$ recovers standard cross-entropy; choosing a uniform distribution $\alpha=(\tfrac{1}{C},\dots,\tfrac{1}{C})$ yields a ListMLE-like objective; and any fixed, strictly decreasing sequence reproduces the P-ListMLE surrogate.
+Furthermore, our data-driven weights recover several known ranking surrogates as special cases.
+Setting $\alpha=(1,0,\dots,0)$ gives the standard cross-entropy.
+A uniform $\alpha=(\tfrac{1}{C},\dots,\tfrac{1}{C})$ produces a ListMLE-like objective, and any fixed decreasing sequence yields the P-ListMLE surrogate.
 (a) Top-1 accuracy in the homogeneous distillation setting.
@@
 In particular, each summand
-$-\,s_{\pi^{*}_{k}}\;+\;\log\!\sum_{\ell=k}^{C}e^{s_{\pi^{*}_{\ell}}}$
+$s_{\pi^{*}_{k}}\;+\;\log\!\sum_{\ell=k}^{C}e^{s_{\pi^{*}_{\ell}}}$
 is convex in the logits $s$, as it is the sum of an affine function and a log-sum-exp.
@@
 5 Experiments
-All experiments are conducted on ImageNet-1K deng2009imagenet following the recipe wightman2021resnet with minimal data augmentation.
+We evaluate PLD on three representative visual recognition datasets: CIFAR-100 cifar100 , ImageNet-1K deng2009imagenet , and MS-COCO coco .
+These datasets cover small- and large-scale image classification and general object detection.
+We first benchmark PLD on CIFAR-100 using traditional convolutional architectures and a unified training recipe.
+We then assess scalability on ImageNet-1K, where we adopt stronger architectures and optimized training strategies.
+Finally, we test PLD on MS-COCO for object detection to examine generalization beyond classification.
+5.1 CIFAR-100 Classification
+We compare PLD with classical and feature-based knowledge distillation methods on CIFAR-100 cifar100 .
+The dataset contains 50,000 training images and 10,000 validation images across 100 categories at $32\times 32$ resolution.
+All models are trained from scratch for 250 epochs.
+We use the AdamW optimizer with $\beta_{1}{=}0.9$ and $\beta_{2}{=}0.999$.
+The learning rate follows a cosine-annealing schedule, starting at 0.001.
+We set weight decay to 0.5 and the batch size to 128.
+Table 1: Evaluation results on the CIFAR-100 dataset. The upper and lower models denote the teacher and student, respectively.
+Method Homogeneous setup Heterogeneous setup
+[…]
+[…]
+Feature-based methods
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+Logits-based methods
+[…]
+[…]
+[…]
+[…]
+We apply standard data augmentations: random cropping, horizontal flipping, and per-channel normalization.
+We report the mean and standard deviation over three independent runs.
+For DIST huang2022knowledge , we set $\alpha{=}1.0$ for the cross-entropy term, $\beta{=}2.0$ for the inter-class relation loss, $\gamma{=}2.0$ for the intra-class relation loss, and use a temperature of $\tau{=}4.0$.
+PLD uses the same temperature ($\tau{=}4.0$).
+5.2 ImageNet-1K Classification
+All experiments on ImageNet-1K deng2009imagenet are conducted following the recipe wightman2021resnet with minimal data augmentation.
 We train for 100 epochs with an effective batch size of 2048 images (256 per GPU across eight NVIDIA A100 SXM4 80 GB accelerators) using the LAMByou2019large optimizer, an initial learning rate of $5\times 10^{-3}$ decayed by a cosine schedule and linearly warmed up over the first 5 epochs, and weight decay fixed at 0.02.
 Under this setup, 100 epochs require approximately 6 hours for convolutional backbones and 8 hours for Vision Transformers.
-To focus on the impact of our loss formulations, we apply only random resized crops, horizontal flips, and per-channel normalization to ImageNet statistics.
-For validation we adopt the “A-recipe” from wightman2021resnet .
+To isolate the effect of our losses, we use only random resized crops, horizontal flips, and per-channel normalization.
+For validation we adopt the "A-recipe" from wightman2021resnet .
 Specifically, we set the test resolution $r=224$ and test crop ratio $\rho=0.95$, then $\text{resize}_{\min}=\bigl\lceil r/\rho\bigr\rceil\approx 236,$ apply a bicubic resize of the shorter side to $\text{resize}_{\min}$, followed by a center crop of size $r\times r$.
-Finally, we normalize by the same mean and standard deviation used during training.
-5.1 Baseline Configuration and Ablation
-We evaluate four losses-CE, KD hinton2015distilling , DIST huang2022knowledge , and PLD-using a ResNet-50 student and ResNet-152 teacher He2015 under identical training.
-Each objective is a drop-in replacement of the logit-based term and incurs comparable cost, so KD (divergence-based) and DIST (correlation-based) serve as natural baselines against our ranking-based PLD.
-We sweep each method’s key hyperparameters ($\alpha$ for KD, $\beta,\gamma$ for DIST, $\tau_{T}$ for PLD) to find optimal settings.
-Figure 1(a) shows the effect of varying the CE mixing weight: gradually reducing $\alpha$ improves both KD and DIST up to an optimal region near $\alpha\approx 0.1$, but fully dropping CE ($\alpha=0$) degrades performance.
-Table 1 reports results with and without the CE term (full ablation in Appendix B), and Table 2 gives PLD-specific sweeps.
-Table 1: Top-1 and Top-5 accuracy for baseline methods (ResNet-50 student, ResNet-152 teacher with Top-1 Acc.
-79.61%).
-“–” indicates not applicable.
-CE: standard cross-entropy.
-LS:szegedy2016rethinking cross-entropy with label smoothing ($\epsilon=0.1$).
-KD: vanilla distillation with mixing weight $\alpha$ and temperature $\tau=2$.
-DIST: relational distillation with inter-class weight $\beta$, intra-class weight $\gamma$, and temperature $\tau=1$.
-Method $\alpha$ $\beta$ $\gamma$ $\tau$ Top-1 Acc.Top-5 Acc.%
-Teacher (ResNet-152) 79.61
-CE 71.35
-LS ($\epsilon=0.1$) 73.92
-KD 0.00 2.00 75.92 92.82
-0.10 2.00 76.80 93.16
-DIST 0.00 0.50 0.50 1.00 76.47 93.19
-0.10 0.45 0.45 1.00 76.60 93.30
-Table 2: Top-1 and Top-5 accuracy for PLD variants (ResNet-50 student, ResNet-152 teacher with Top-1 Acc. 79.61%). “–” indicates not applicable.
-Variant $\tau$ Top-1 Acc.Top-5 Acc.%
-PLD 0.50 76.06 92.39
-1.00 77.30 93.28
-1.50 77.17 92.74
-2.00 76.46 92.34
-4.00 75.22 91.34
-PLD–ListMLE 74.11 90.60
-PLD–pListMLE 76.60 93.23
-5.1.1 Ablation of PLD
-As PLD weights each position by $\alpha_{k}\;=\;\frac{\exp\bigl{(}t_{\pi^{*}_{k}}/\tau_{T}\bigr{)}}{\sum_{j=1}^{C}\exp\bigl{(}t_{\pi^{*}_{j}}/\tau_{T}\bigr{)}}$, we sweep $\tau_{T}\in\{0.5,1.0,1.5,2.0,4.0\}$ to assess sensitivity: lower $\tau_{T}$ sharpens the distribution, while higher $\tau_{T}$ flattens it.
-We also test two special cases: Uniform (ListMLE): $\alpha_{k}=1/C$.
-Position-aware (p-ListMLE): $\alpha_{k}=q^{T}_{\pi^{*}_{k}}/\sum_{j}q^{T}_{\pi^{*}_{j}}$.
-As Table 2 shows, the unsoftened PLD ($\tau_{T}=1$) consistently achieves the highest Top-1/Top-5 accuracies, with uniform ListMLE trailing by 2-3 pp.
-Overall, the best results occur at $\tau_{T}=1.0$ (no softening) or with only slight perturbations ($1.0\pm\epsilon$), demonstrating that PLD is robust to the exact choice of $\alpha$ distribution and does not require aggressive sharpening or over-softening.
-5.2 Key Experiments
-To evaluate PLD’s robustness and generality, we conduct three complementary studies using hyperparameters identified in our baseline experiments (KD: $\alpha=0.1,\tau=2$; DIST: $\beta=0.45,\gamma=0.45,\tau=1$; PLD: $\tau_{T}=1$).
-First, in a homogeneous setting, we pair larger teachers with smaller students within the same architecture family to isolate the impact of model capacity on distillation.
-Second, in a heterogeneous setup, we fix a single student (ResNet-50) and vary teacher architectures across convolutional and transformer families to test PLD under diverse teacher distributions.
-Finally, we return to the homogeneous scenario with extended pretraining, pretraining a MobileNet-v4qin2024mobilenetv4 Medium student from a MobileNet-v4 Large teacher for 300 epochs to verify that PLD’s gains persist under longer optimization.
-5.2.1 Distillation Across Homogeneous Architectures
-As illustrated in Figure 2(a), we evaluate three backbone families-ResNetHe2015 , MobileNet-v4qin2024mobilenetv4 , and Vision Transformersdosovitskiy2020image -in a homogeneous setting.
-Across convolutional models (ResNet, MobileNet-v4), PLD outperforms KD by 0.5-2.7 pp and DIST by 0.2-0.7 pp.
-On Vision Transformers (ViT), PLD yields +0.72 pp over DIST and +0.30 pp over KD.
-Averaged across all pairs, PLD improves Top-1 accuracy by +0.42 pp versus DIST and +1.04 pp versus KD.
-Table 3 reports the full Top-1 results for KD, DIST, and PLD.
-5.2.2 Distillation Across Diverse Teachers
-One of the key benefits of logit-based distillation is its ability to leverage teacher models of different architectures, provided their output logits share the same dimensionality as the student.
-Figure 2(b) shows the heterogeneous setting: we fix the student to ResNet-50 ($\approx 25.56$ M parameters) and distill from a range of convolutional and transformer teachers.
-Across all teachers, PLD consistently outperforms both KD and DIST: its $\Delta_{\mathrm{DIST}}$ gains range from +0.34 pp (MobileNet-v4 Hybrid) to +0.70 pp (ResNet-152), averaging +0.48 pp, while $\Delta_{\mathrm{KD}}$ gains span +0.49 pp to +1.55 pp, with an average of +1.09 pp.
-As teacher capacity increases, KD’s benefit declines and DIST shows a modest upward trend; PLD not only follows this trend but further widens the margin over DIST for larger teachers.
-The largest gap between DIST and PLD appears in the homogeneous ResNet-50←ResNet-152 pairing.
-Table 4 reports each teacher’s standalone Top-1 accuracy, parameter count, and student Top-1 results under KD, DIST, and PLD, along with $\Delta_{\mathrm{DIST}}$ and $\Delta_{\mathrm{KD}}$.
+We normalize by the same mean and standard deviation as in training.
 Table 2: Top-1 accuracy (%) and student model size across student–teacher pairs on ImageNet-1K.
-“Teacher Acc” is the teacher’s standalone Top-1; $\Delta_{\mathrm{DIST}}$ and $\Delta_{\mathrm{KD}}$ are PLD’s gains over DIST and KD.
+"Teacher Acc" is the teacher’s standalone Top-1; $\Delta_{\mathrm{DIST}}$ and $\Delta_{\mathrm{KD}}$ are PLD’s gains over DIST and KD.
 Top-1 % Accuracy
@@
 Table 3: Top-1 % accuracy of ResNet-50 distilled from diverse teachers on ImageNet-1K.
-“Params” is the teacher’s parameter count in millions.
+"Params" is the teacher’s parameter count in millions.
 $\Delta_{\mathrm{DIST}}$ and $\Delta_{\mathrm{KD}}$ report PLD’s gains over DIST and KD, respectively.
@@
 ViT-Large/16 304.33 84.80 75.98 76.86 77.38 0.52 1.40
-5.2.3 Consistency under Extended Training
-As observed in our homogeneous experiments5.2.1 -where KD sometimes outperforms DIST and vice versa-we sought to confirm whether PLD consistently surpasses both.
-To this end, we independently trained a MobileNet-v4 Medium student from a MobileNet-v4 Large teacher for 100 and 300 epochs using identical hyperparameters.
-Table 5 reports Top-1 accuracies on ImageNet-1K, and Figure 1(b) provides a detailed comparison.
-After 100 epochs, PLD achieves 75.72% (vs.
-DIST +0.32 pp, KD +0.86 pp), rising to 76.94% at 300 epochs (vs.
-DIST +0.38 pp, KD +0.80 pp).
-We observe that the distillation gains with PLD (+1.22 pp) are nearly twice those of standard pretraining (+0.67 pp), and that this scaling is comparable across KD (+1.28 pp) and DIST (+1.16 pp).
-This demonstrates that PLD not only preserves but slightly widens its distillation advantage under extended training.
 Table 4: Top-1 accuracy (%) of the MobileNet-v4 Medium student distilled from a MobileNet-v4 Large teacher after 100 and 300 training epochs.
-The “$\Delta$” row shows the improvement from 100 to 300 epochs.
+The $\Delta$" row shows the improvement from 100 to 300 epochs.
 Epochs CE KD DIST PLD $\Delta_{\mathrm{DIST}}$ $\Delta_{\mathrm{KD}}$
@@
 $\Delta$ (300–100) 0.67 1.28 1.16 1.22
+5.2.1 Baseline Configuration and Ablation
+We evaluate four losses under identical training: CE, KD hinton2015distilling , DIST huang2022knowledge , and PLD.
+We use a ResNet-50 student and a ResNet-152 teacher He2015 .
+Each objective replaces only the logit-based term and has comparable cost.
+Thus, KD (divergence-based) and DIST (correlation-based) are natural baselines for our ranking-based PLD.
+We sweep the key hyperparameters for each method: $\alpha$ for KD, $\beta,\gamma$ for DIST, and $\tau_{T}$ for PLD.
+Figure 1(a) shows the effect of the CE mixing weight.
+Reducing $\alpha$ improves KD and DIST up to a region near $\alpha\approx 0.1$.
+Setting $\alpha=0$ (no CE) degrades performance.
+Table 5 reports results with and without CE (see Table 8 for the full ablation).
+Table 6 lists PLD-specific sweeps.
+5.2.2 Ablation of PLD
+As PLD weights each position by $\alpha_{k}\;=\;\frac{\exp\bigl(t_{\pi^{*}_{k}}/\tau_{T}\bigr)}{\sum_{j=1}^{C}\exp\bigl(t_{\pi^{*}_{j}}/\tau_{T}\bigr)}$, we sweep $\tau_{T}\in\{0.5,1.0,1.5,2.0,4.0\}$ to assess sensitivity: lower $\tau_{T}$ sharpens the distribution, while higher $\tau_{T}$ flattens the distribution.
+We also test two special cases: Uniform (ListMLE): $\alpha_{k}=1/C$.
+Position-aware (p-ListMLE): $\alpha_{k}=q^{T}_{\pi^{*}_{k}}/\sum_{j}q^{T}_{\pi^{*}_{j}}$.
+As Table 6 shows, the unsoftened PLD ($\tau_{T}=1$) consistently achieves the highest Top-1/Top-5 accuracies, with uniform ListMLE trailing by 2-3 pp.
+Overall, the best results occur at $\tau_{T}=1.0$ (no softening) or with only slight perturbations ($1.0\pm\epsilon$), demonstrating that PLD is robust to the exact choice of $\alpha$ distribution and does not require aggressive sharpening or over-softening.
+Table 5: Top‐1 and Top‐5 accuracy for baseline methods (ResNet-50 student, ResNet-152 teacher with Top-1 Acc.
+79.61%).
+"–" indicates not applicable.
+CE: standard cross‐entropy.
+LS:szegedy2016rethinking cross‐entropy with label smoothing ($\epsilon=0.1$).
+KD: vanilla distillation with mixing weight $\alpha$ and temperature $\tau=2$.
+DIST: relational distillation with inter-class weight $\beta$, intra-class weight $\gamma$, and temperature $\tau=1$.
+Method $\alpha$ $\beta$ $\gamma$ $\tau$ Top-1 Acc.% Top-5 Acc.%
+Teacher (ResNet-152) 79.61
+CE 71.35
+LS ($\epsilon=0.1$) 73.92
+KD 0.00 2.00 75.92 92.82
+0.10 2.00 76.80 93.16
+DIST 0.00 0.50 0.50 1.00 76.47 93.19
+0.10 0.45 0.45 1.00 76.60 93.30
+Table 6: Top‐1 and Top‐5 accuracy for PLD variants (ResNet‐50 student, ResNet‐152 teacher with Top‐1 Acc. 79.61%). "–" indicates not applicable.
+Variant $\tau$ Top-1 Acc.% Top-5 Acc.%
+PLD 0.50 76.06 92.39
+1.00 77.30 93.28
+1.50 77.17 92.74
+2.00 76.46 92.34
+4.00 75.22 91.34
+PLD–ListMLE 74.11 90.60
+PLD–pListMLE 76.60 93.23
+5.2.3 Distillation Across Homogeneous Architectures
+Figure 2(a) evaluates three backbone families in a homogeneous setting: ResNet He2015 , MobileNet-v4 qin2024mobilenetv4 , and Vision Transformers dosovitskiy2020image .
+Across convolutional models, PLD outperforms KD by 0.5–2.7 pp and DIST by 0.2–0.7 pp.
+On ViT, PLD yields +0.72 pp over DIST and +0.30 pp over KD.
+Averaged over all pairs, PLD improves Top-1 accuracy by +0.42 pp versus DIST and +1.04 pp versus KD.
+Table 2 reports full Top-1 results for KD, DIST, and PLD.
+5.2.4 Distillation Across Diverse Teachers
+[…]
+5.2.5 Consistency under Extended Training
+Our homogeneous experiments (Sec.
+5.2.3) show that KD sometimes beats DIST and vice versa.
+We ask whether PLD consistently surpasses both under longer training.
+We train a MobileNet-v4 Medium student from a MobileNet-v4 Large teacher for 100 and 300 epochs with identical hyperparameters.
+Table 4 reports Top-1 on ImageNet-1K, and Figure 1(b) provides a detailed comparison.
+At 100 epochs, PLD achieves 75.72%.
+This is +0.32 pp over DIST and +0.86 pp over KD.
+At 300 epochs, PLD reaches 76.94% (+0.38 pp over DIST; +0.80 pp over KD).
+The PLD gains with longer training are +1.22 pp, which is nearly twice the standard pretraining gains (+0.67 pp).
+This scaling is comparable across KD (+1.28 pp) and DIST (+1.16 pp).
+Thus, PLD preserves and slightly widens its advantage under extended training.
+Table 7: Object detection performance on MS-COCO.
+We distill Faster R-CNN detectors with FPN lin2017feature using DKD zhao2022decoupled and PLD.
+All models are trained under identical schedules.
+Results are reported on val2017.
+Teacher $\rightarrow$ Student Method AP AP50 AP75 APs APm APl
+ResNet-50 $\rightarrow$ MobileNetV2-FPN DKD 29.24 50.42 29.87 16.15 31.11 38.07
+ResNet-50 $\rightarrow$ MobileNetV2-FPN PLD 28.91 49.63 29.88 16.05 30.41 38.59
+ResNet-101 $\rightarrow$ ResNet-18-FPN DKD 32.11 53.43 33.90 18.46 34.43 41.50
+ResNet-101 $\rightarrow$ ResNet-18-FPN PLD 32.47 53.83 34.17 18.50 34.97 42.12
+ResNet-101 $\rightarrow$ ResNet-50-FPN DKD 36.54 58.57 39.46 21.74 39.80 47.31
+ResNet-101 $\rightarrow$ ResNet-50-FPN PLD 36.60 58.28 39.58 21.37 39.76 47.24
+5.3 Object Detection on MS-COCO
+PLD extends to tasks where the teacher and student share the same output logits, including object detection and semantic segmentation.
+We distill Faster R-CNN lin2017feature detectors on MS-COCO coco with standard FPNs.
+We evaluate two convolutional teacher–student pairs: ResNet-50 $\rightarrow$ MobileNetV2-FPN and ResNet-101 $\rightarrow$ {ResNet-18-FPN, ResNet-50-FPN}.
+All models use identical hyperparameters, and schedules.
+We compare PLD to the strong DKD baseline zhao2022decoupled .
+Table 7 shows that PLD achieves comparable or slightly better performance than DKD across multiple pairs.
+PLD improves AP and AP75 in most settings while maintaining similar APs for small objects.
+These results indicate that PLD transfers structured knowledge for dense prediction without modifying the detector.
 6 Conclusion and Limitation
-In this work, we introduced Plackett-Luce Distillation (PLD), a unified choice-theoretic framework for logit-based knowledge distillation.
-Empirically, PLD consistently outperforms both classical KL-based distillation (KD) and recent correlation-based methods (DIST) across a wide range of student-teacher pairs-homogeneous and heterogeneous architectures alike-and remains robust under extended training.
-These results demonstrate that transferring the teacher’s structured preferences yields stronger and more stable student performance than marginal- or correlation-based objectives.
-While PLD applies broadly to any setting in which student and teacher share the same logit dimensionality, it assumes fully aligned class vocabularies and does not directly support mismatched label sets or incremental class addition.
-Moreover, PLD requires a sorting operation over the $C$ output logits to extract the teacher-optimal permutation, incurring an $O(C\log C)$ complexity per example; although efficient in distributed implementations, this contrasts with the $O(C)$ cost of standard KD or DIST.
-Finally, as with other distillation methods, PLD’s benefits scale with teacher confidence and may diminish when the teacher’s softmax distribution is nearly uniform.
-PLD’s choice-theoretic foundation suggests several promising directions.
-First, adaptive or curriculum-driven weighting schemes could modulate the ranking loss according to sample difficulty.
-Second, extending PLD to other domains-such as sequence modeling, reinforcement learning, or multitask settings-may yield similar gains.
-We hope this work inspires continued exploration of ranking-based objectives for principled and effective model compression.
+In this work, we introduced Plackett–Luce Distillation (PLD), a unified, choice-theoretic framework for logit-based knowledge distillation.
+Empirically, across CIFAR-100, ImageNet-1K, and MS-COCO, PLD achieves consistent gains across diverse architectures and distillation objectives, including divergence-based, correlation-based, and feature-based methods, in both homogeneous and heterogeneous teacher–student pairs.
+These results show that transferring the teacher’s structured preferences yields stronger and more stable student performance than marginal- or correlation-based objectives.
+PLD applies broadly to any setting where the student and teacher share the same logit dimensionality.
+However, it assumes fully aligned class vocabularies and does not yet support mismatched label sets or incremental class addition.
+PLD also requires sorting over the $C$ output logits to extract the teacher-optimal permutation, which has $O(C\log C)$ complexity per example.
+Although efficient in distributed implementations, this cost is higher than the $O(C)$ complexity of standard KD or DIST.
+Empirical runtime analysis (Appendix I) shows that the overhead is negligible for common architectures.
+As with other distillation methods, PLD’s benefits depend on teacher confidence and may diminish when the teacher’s softmax distribution is nearly uniform, as illustrated in the loss landscape analysis (Appendix J).
+PLD’s choice-theoretic foundation opens several directions for future work.
+First, adaptive or curriculum-driven weighting schemes could adjust the ranking loss based on sample difficulty.
+Second, extending PLD to other domains, such as sequence modeling, reinforcement learning, or multitask learning, may offer similar gains.
+We hope this work encourages further exploration of ranking-based objectives for principled and effective model compression.
+Acknowledgments and Disclosure of Funding
+This work is partially sponsored by National Key R&D Program of China under Grant 2022YFB3104200, NSFC U24A20235 and 62032003.
 Appendix A Gradient Derivation for PLD Loss
@@
 $\phi_{k}(s)=\log\sum_{\ell=k}^{C}\exp(s_{\pi^{*}_{\ell}}).$
-Our goal is to compute the gradient s($\nabla_{s}\mathcal{L}(s)$, whose $i$-th component is $\frac{\partial\mathcal{L}}{\partial s_{i}}$. Due to the linearity of differentiation,
+Our goal is to compute the gradient $\nabla_{s}\mathcal{L}(s)$, whose $i$-th component is $\frac{\partial\mathcal{L}}{\partial s_{i}}$. Due to the linearity of differentiation,
 $\frac{\partial\mathcal{L}}{\partial s_{i}}=\sum_{k=1}^{C}\frac{\partial L_{k}}{\partial s_{i}}.$
@@
 1. Derivative of the affine term:
-The first term is $-\alpha_{k}s_{\pi^{*}_{k}}$.
+The first term is $\alpha_{k}s_{\pi^{*}_{k}}$.
 $\frac{\partial}{\partial s_{i}}\left(-\alpha_{k}s_{\pi^{*}_{k}}\right)=-\alpha_{k}\frac{\partial s_{\pi^{*}_{k}}}{\partial s_{i}}.$
@@
 $=\alpha_{k}\left[\sigma_{k}(i)-\mathbf{1}\{i=\pi^{*}_{k}\}\right].$
-A.3 Final Gradient s($\nabla_{s}\mathcal{L}(s)$:
+A.3 Final Gradient $\nabla_{s}\mathcal{L}(s)$:
 The $i$-th component of the gradient of the total loss $\mathcal{L}(s)$ is:
-$\frac{\partial\mathcal{L}}{\partial s_{i}}$ $=\sum_{k=1}^{C}\frac{\partial L_{k}}{\partial s_{i}}$
-$=\sum_{k=1}^{C}\alpha_{k}\left[\sigma_{k}(i)-\mathbf{1}\{i=\pi^{*}_{k}\}\right].$
+$\frac{\partial\mathcal{L}}{\partial s_{i}}$ $=\sum_{k=1}^{C}\frac{\partial L_{k}}{\partial s_{i}}$ $=\sum_{k=1}^{C}\alpha_{k}\left[\sigma_{k}(i)-\mathbf{1}\{i=\pi^{*}_{k}\}\right].$
 We can separate this sum into two parts:
@@
 A.4 Convexity and Smoothness:
-Each term $L_{k}(s)=\alpha_{k}\left[-s_{\pi^{*}_{k}}+\log\sum_{\ell=k}^{C}\exp(s_{\pi^{*}_{\ell}})\right]$ is a sum of an affine function ($-\alpha_{k}s_{\pi^{*}_{k}}$) and a non-negatively weighted log-sum-exp function ($\alpha_{k}\phi_{k}(s)$).
+Each term $L_{k}(s)=\alpha_{k}\left[-s_{\pi^{*}_{k}}+\log\sum_{\ell=k}^{C}\exp(s_{\pi^{*}_{\ell}})\right]$ is a sum of an affine function ($\alpha_{k}s_{\pi^{*}_{k}}$) and a non-negatively weighted log-sum-exp function ($\alpha_{k}\phi_{k}(s)$).
 Affine functions are convex.
@@
 Appendix B Full Baseline Ablation
-Table 6: Top-1 and Top-5 accuracy for baseline methods (ResNet-50 student, ResNet-152 teacher with Top-1 Acc.
+Table 8: Hyperparameter grid search on ImageNet-1K.
+Each configuration uses a ResNet-50 student distilled from a ResNet-152 teacher (Top-1 Acc.
 79.61%).
-“-” indicates not applicable.
-CE: standard cross-entropy.
-LS: cross-entropy with label smoothing ($\epsilon=0.1$).
-KD: vanilla distillation with mixing weight $\alpha$ and temperature $\tau=2$.
-DIST: relational distillation with inter-class weight $\beta$, intra-class weight $\gamma$, and temperature $\tau=1$.
-Method $\alpha$ $\beta$ $\gamma$ $\tau$ Top-1 Acc.Top-5 Acc.%
-Oracle Teacher
-Teacher (ResNet-152) - - - - 79.61 -
-Baselines
-CE - - - - 71.35 -
-LS ($\epsilon=0.1$) - - - - 73.92 -
-Vanilla KD ($\tau=2$)
-KD 0.00 - - 2.00 75.92 92.82
-0.10 - - 2.00 76.80 93.16
-0.20 - - 2.00 76.65 93.11
-0.30 - - 2.00 76.30 93.02
-0.40 - - 2.00 76.33 93.10
-0.50 - - 2.00 76.12 93.12
-0.60 - - 2.00 75.94 92.79
-0.70 - - 2.00 75.85 92.71
-0.80 - - 2.00 75.21 92.39
-0.90 - - 2.00 74.28 91.59
-DIST ($\tau=1$)
-DIST 1.00 2.00 2.00 1.00 75.77 92.67
+"–" indicates not applicable.
+The best setting in each group is highlighted.
+Method Hyperparameters Accuracy (%)
+$\alpha$ $\beta$ $\gamma$ $\tau$ Top-1 Top-5
+Teacher 79.61
+CE 71.35
+LS ($\epsilon{=}0.1$) 73.92
+KD hyperparameter sweep ($\tau{=}2$)
+0.00 2.00 75.92 92.82
+KD 0.10 2.00 76.80 93.16
+0.20 2.00 76.65 93.11
+0.30 2.00 76.30 93.02
+0.40 2.00 76.33 93.10
+0.50 2.00 76.12 93.12
+0.60 2.00 75.94 92.79
+0.70 2.00 75.85 92.71
+0.80 2.00 75.21 92.39
+0.90 2.00 74.28 91.59
+DIST hyperparameter sweep ($\tau{=}1$)
+1.00 2.00 2.00 1.00 75.77 92.67
 0.00 0.50 0.50 1.00 76.47 93.19
-0.10 0.45 0.45 1.00 76.60 93.30
+DIST 0.10 0.45 0.45 1.00 76.60 93.30
 0.20 0.40 0.40 1.00 75.69 92.63
@@
 To ensure that any observed differences arise solely from the choice of loss function, we use exactly the same training hyperparameters as in our main experiments, including learning-rate schedules, batch size, weight decay, and the data-augmentation pipeline.
-To comprehensively evaluate the relative merits of Plackett-Luce Distillation (PLD) compared to DIST and KD, we conduct experiments across multiple optimizers, various divergence weightings, and several logit-standardization schemes.
+To comprehensively evaluate the relative merits of Plackett-Luce Distillation (PLD) compared to DIST and KD, we conduct experiments across multiple optimizers, various divergence weightings, runtime and several logit-standardization schemes.
 We further quantify distributional alignment by measuring the KL divergence between student and teacher softmax outputs throughout training.
@@
 ViT-Large/16 ResNet50 0.46 0.44 0.59
-Appendix I Loss Landscape Visualization
+Table 13: Total training time (minutes) for all CIFAR-100 experiments. PLD achieves comparable efficiency to KD, DIST, and DKD despite its list-wise formulation.
+Model (Teacher Student) CE PLD KD DIST DKD AT FitNet PKT RKD SP VID
+[…]
+ResNet50 MobileNetV2 23.62 102.59 102.23 102.49 102.54 108.17 113.18 102.28 128.65 102.78 132.06
+ResNet56 ResNet20 22.97 33.42 33.57 33.81 33.46 34.55 33.30 33.89 36.21 33.67 36.57
+WideResNet-40-2 WideResNet-40-1 27.67 33.10 37.56 36.73 34.27 36.77 38.05 36.84 37.62 37.01 37.35
+Appendix I Runtime Analysis
+We evaluate the computational efficiency of PLD compared with other distillation methods discussed in Section 5.
+PLD involves a single sorting operation to obtain the teacher-optimal permutation, similar to ListMLE xia2008listwise ; therefore, its complexity is only $O(C\log C)$ far lower than the $O(C!)$ enumeration of all possible rankings.
+In practice, this adds negligible overhead: PLD trains at nearly the same speed as logit-based methods such as KD, DIST, and DKD.
+Table 13 reports the total training time (in minutes) for all CIFAR-100 experiments in Table 1.
+Despite its list-wise formulation, PLD’s runtime remains comparable to KD and DIST, and substantially faster than other methods such as RKD or VID.
+Appendix J Loss Landscape Visualization
+(a) $T=2.0$
+(b) $T=1.0$
+(c) $T=0.5$
+(d) $T=0.1$
+Figure 3: PLD loss surfaces at different teacher temperatures. (Top row) $T=2.0$ and $T=1.0$; (Bottom row) $T=0.5$ and $T=0.1$. Lowering $T$ below 1.0 flattens convexity.
 To gain geometric insight into the optimization landscapes induced by our three distillation losses (DIST, KD, PLD), we plot 3D surfaces and 2D contours over a two-dimensional slice of the student-logit space.
@@
 Figure 4: Loss landscapes of three distillation methods: (a) DIST exhibits a sharp dip yet remains effectively planar; (b) KD shows moderate convexity; (c) PLD (ours) exhibits better convexity with contours mostly centered at the origin.
-I.1 Temperature Sensitivity of the PLD Loss Surface
+J.1 Temperature Sensitivity of the PLD Loss Surface
 To investigate the effect of the teacher-softmax temperature $T$ on the geometry of the PLD loss landscape, we fix the same two-dimensional $(\alpha,\beta)$ slice and compute the PLD surface at four representative temperatures: $T\in\{2.0,1.0,0.5,0.1\}$.
@@
 We observe that reducing $T$ from 2.0 down to 1.0 produces only minor changes, whereas further lowering $T$ below 1.0 flattens the curvature.
-(a) $T=2.0$
-(b) $T=1.0$
-(c) $T=0.5$
-(d) $T=0.1$
-Figure 4: PLD loss surfaces at different teacher temperatures. (Top row) $T=2.0$ and $T=1.0$; (Bottom row) $T=0.5$ and $T=0.1$. Lowering $T$ below 1.0 flattens convexity.
+NeurIPS Paper Checklist
+  1. 1.
+    Claims
+  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
+  3. Answer: [Yes]
+  4. 2.
+    Limitations
+  5. Question: Does the paper discuss the limitations of the work performed by the authors?
+  6. Answer: [Yes]
+  7. 3.
+    Theory assumptions and proofs
+  8. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
+  9. Answer: [Yes]
+  10. 4.
+    Experimental result reproducibility
+  11.
+Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
+  12. Answer: [Yes]
+  13. 5.
+    Open access to data and code
+  14. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
+  15. Answer: [Yes]
+  16. 6.
+    Experimental setting/details
+  17. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
+  18. Answer: [Yes]
+  19. 7.
+    Experiment statistical significance
+  20. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
+  21. Answer: [Yes]
+  22. Justification: We report mean Top-1 accuracies on CIFAR-100 along with their corresponding standard errors computed over three independent runs with different random seeds.
+  23. 8.
+    Experiments compute resources
+  24. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
+  25. Answer: [Yes]
+  26. 9.
+    Code of ethics
+  27. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
+  28. Answer: [Yes]
+  29. 10.
+    Broader impacts
+  30. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
+  31. Answer: [N/A]
+  32.
+Justification: This work is foundational on model compression and does not present a direct application or risk.
+Indirectly, more efficient models may reduce energy and hardware costs, but we do not see specific societal harms arising from the method itself.
+  33. 11.
+    Safeguards
+  34.
+Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
+  35. Answer: [N/A]
+  36. Justification: No high-risk or dual-use data/models are released in this work
+  37. 12.
+    Licenses for existing assets
+  38.
+Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
+  39. Answer: [Yes]
+  40. 13.
+    New assets
+  41. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
+  42. Answer: [N/A]
+  43. Justification: We do not release any new datasets or large models with this submission.
+  44. 14.
+    Crowdsourcing and research with human subjects
+  45.
+Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
+  46. Answer: [N/A]
+  47. Justification: No human-subject data or crowdsourcing is involved.
+  48. 15.
+    Institutional review board (IRB) approvals or equivalent for research with human subjects
+  49.
+Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
+  50. Answer: [N/A]
+  51. Justification: This work does not involve human subjects.
+  52. 16.
+    Declaration of LLM usage
+  53.
+Question: Does the paper describe the usage of LLMs if it is an important, original, or non-standard component of the core methods in this research?
+Note that if the LLM is used only for writing, editing, or formatting purposes and does not impact the core methodology, scientific rigorousness, or originality of the research, declaration is not required.
+  54. Answer: [N/A]
+  55. Justification: No large language models are used in our core methodology.